A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.01

Title word cross-reference

0-Year [KH55].

1950 [Ano50j]. 1951 [Ano51f]. 1955 [Ano55f]. 1958 [Ano58g, Ano59c].

Ability [KPW59]. Abundance [CLRC53, Cro55, Glu55, SK54]. Acid [Wat57]. Acipenser [PC55]. AD [Ano52i]. Added [And50, Bar53]. Administrative [Wol52]. Adsorption [Bar53]. Advancements [All54, Ste58]. Aerial [HB58]. Affected [LCH59]. Age [Alv54, AS51, BS50, Ket50, LaP58, McC52, McF59, Neu57, PC55, Roe58, Sig52]. Agents [WD59]. Aging [Hav59]. Aid [Car58b]. Alabama [BM57, Sco51, Swi54, Whi59]. Alaska [Roo59]. Alberta [Mil52, MT57a, RE50]. Alewife [Mil57]. Algonquin [Woo53, Lan50, Mar52, Mar57]. Alkalinity [Bar53]. Alma [Pfl50]. Alosa [Mas52, Mil57, Wal57]. Alternating [Pra52]. America [Nig54]. American [Ano50g, Ano50h, Ano51e, Ano52d, Ano53f, Ano54g, Ano55e, Ano57-27, Ano57i, Ano58f, Ano59c, Ano52a, Ano53c, Ano54c, Ano57e, Ano58c, Bal50a, Hal54a, Hil54, LaP58, Rup59, SSS59, Wal57]. Analysis [WL58]. Angler [MT57b]. Angling [RE50]. Annual [Ano50v, Ano50w, Ano51b, Ano52a, Ano53d, Ano54d, Ano55c, Ano57-27, Ano57f, Ano58d, Ano59c, Ano59g]. Annuli [Spr54b]. Annulus [McF59]. Apatemon [Hof59]. Aplodinotus [BS50]. Appalachians [KH52]. Apparent [And50]. Appearance [BW58]. Application [Joe53, San57, Wat57]. Applied [RB50]. Appointment [Ano50a, Ano50b, Ano51a, Ano52b, Ano53a, Ano54a, Ano55a, Ano57a, Ano57b, Ano58a]. Approach [Ros51a, Tho59]. April [Ano50j, Ano58g]. Aquaria [Wil59]. Aquatic [Ide57, Jac50, VAK50]. Area [Wal51]. Areas [LP55, Mas52]. arenaria [Glu55]. Arizona [Wal51]. Arm [Wit57]. Artificial [Dym57, Eis58, HBT53, HM57, Pre57]. Aspects [Sch59, Whi57]. Assay [HW50, Sai54]. Associated [Gre58, HC51]. Atlantic [Hav59, LaP58, McC50, McH59, RB50, Spr57]. atraria [Neu57]. Attacks [Roy50]. Attempt [RE50]. Attendance [Ano53j, Ano54k, Ano55j, Ano57p, Ano57q, Ano58j, Ano59g]. Atypical [Nig54]. August [Ano55f]. Automatic [WSV57]. Award [Cal54].

Backwater [Lam59]. Bacteria [Gri54]. Bacterial [Sni54]. Bacterium [GSF53]. bairdi [Zar52]. Balance [Car58a]. Based [Car50b]. Basin [McL52]. Basis [Wol52]. Bass [Ben51, Bon53, Bro51, COHB57, CLRC53, Dav59, DH50, FW57, Gri59, HF57, LE53, MT57b, Par54, PC59, Pra51, Sai58, Scr57, Swi51, Syk58, Wil59]. Basses [Vio53]. Bay [Bal50a, Fry53, Hal54a, Mra52, Roe58, Tha59, Bud57, Joe57]. Beaver [Rup55]. Beaver-Trout [Rup55]. becomes [Car52]. Beds [Pre57]. below [Par57]. Between [Bar53, Ech55, Hun53, Bee58, Kat51a]. Bibliographic [Sca54]. bicinctus [Sur59]. Big [Roe58]. Bio [Sai54]. Bio-Assay [Sai54]. Biological [CF57, Sch59]. Biologists [Sca54]. Biology [Hac57, Hue59, Idy58, MD59]. Black [Bro51, HF57, Hui54, Rei50, Sai58, Vio53]. Blindness [All51]. Blood [Kat51b]. Blue [Dar59, Wol57]. Blue-Sac [Wol57]. Bluegill [Byr52]. Bluegills [Spr54b, Swi51]. Bog [Wat57]. Bone [McC52]. Bones [Neu57]. Book [Ano59a, Ano59b]. Bottom [Bri50, CO57, Rug59, Woo53]. Breeding [DO57]. Brine [CJ55, CF57]. Brine-Polluted [CF57]. Britain [Wol52]. British [CL59, LS54, ML55]. Brood [All59, DO57]. Brook [All51, Alv54, Bal57, Ben54, Coo52a, Coo53, Gre52, HE51, McF59, She50, SGF50, SW55, WYS55]. Brown [Alv54, Coo52a, Coo53, Fla50, Sig52, SW55]. Browns [MT57b]. Burbot [Cle51a, Cle51b]. Business [Ano50c, Ano50d, Ano52c, Ano53b, Ano54b, Ano55b, Ano57c, Ano57d, Ano58b]. By-Laws [Ano58c].

Calculating [Sne51]. California [Bri50, Kim58, MN52, WC55]. Callinectes [Dar59]. Can [McH59]. canadense [Car50a, Car50a]. Canadians [Ken57]. Canal [Hun53]. Candida [ST57]. Capture [Wol51]. Careers [Swi59]. Carolina [Scr57]. Carp [Att59, McC52, Shi58a, AB52]. carpio [McC52]. Cascade [Kat51a]. Catch [Cop57, HG53, MKB50]. Catfish [AS51, Bro51, MD59, Sne51]. Causative [All51]. Cause [BL50]. Caused [Pra55]. Cell [Nig54]. Cells [Saw59]. Celluloid [McK50]. Census [MF57, Mur55, TH51]. Central [Zel53]. Centrarchid [KPW59]. Century [HLN54]. Cepedianum [Mil57]. cephalus [IS52]. cerevisiae [ST57]. Certain [Saw59, Whi59]. Chain [Gri59]. Change [CL59]. Changes [Bec50, Mof57, SK54]. Channel [AS51, Bro51, HM57, Sne51]. Channels [SV59]. Characteristics [LP55, Mas52]. Charts [RB50]. Chemical [LP55]. Chemicals [Saw59]. Chignik [Roo59]. Chilliwack [ML55]. Chinook [Ell58, OF57, Sey59, WC55]. Chipola [PC59]. Chlorides [And50]. Chlorinated [HPT59]. chrysops [Bon53]. Chub [Neu57]. Chubs [KH55]. Circulation [HBT53]. Cisco [Mil50]. Cismontanus [Laa51]. Clam [Glu55]. clarki [Fle52]. clarkii [DHB57]. Class [Fry53, Joh57, KH55]. Classes [FW57]. Claytor [Ros51b]. Clear [CLRC53, McC59]. Clipped [Ger58]. clupeaformis [Roe58]. Coast [IS52, Kna50, LaP58]. Coastal [McH59]. Cold [DB57]. Colorado [Kim58, CO57]. Columbia [Lay50, CL59, HF57, JP59, LS54, ML55, OF57]. Commercial [Ano55i, Bel58, Gre58, McK59, Pow58, Rig58, Sch59, SK54, Whi59]. Commission [Far57]. Committee [Ano53k]. Committees [Ano50a, Ano50b, Ano50k, Ano50l, Ano50t, Ano50r, Ano50p, Ano50u, Ano50s, Ano50q, Ano50z, Ano50x, Ano50-27, Ano50y, Ano51a, Ano51g, Ano51k, Ano51j, Ano51l, Ano52b, Ano52f, Ano52i, Ano52l, Ano52k, Ano52m, Ano53a, Ano53h, Ano53n, Ano53m, Ano53o, Ano54a, Ano54i, Ano54l, Ano54m, Ano55a, Ano55g, Ano55n, Ano55m, Ano55l, Ano55o, Ano57a, Ano57b, Ano57c, Ano57k, Ano57l, Ano57x, Ano57w, Ano57v, Ano57t, Ano57s, Ano57y, Ano57z, Ano58a, Ano58b, Ano58h, Ano58l, Ano58m, Ano59h, Her54, Kru54]. Comparative [Mil54]. Comparison [ST57, Wol52]. Compiled [Ano50j, Ano51f, Ano55f, Ano57j, Ano58g]. Comprehensive [GSF53]. Computation [Hil50]. Concentrations [Zel53]. Concluding [Ano50c, Ano50d, Ano52c, Ano53b, Ano54b, Ano55b, Ano57d]. Condition [CLRC53]. Conditions [CF57, MN52]. Conservation [Ano55i, Kna50]. Consideration [Gab50]. Considerations [Car50b, Ros51a]. Constant [COHB57]. Constant-Temperature [COHB57]. Constitution [Ano53c, Ano54c, Ano57e, Ano58c]. Consumption [Bal57]. Content [KPW59]. Contrast [Hel59]. Contribution [Idy58, SF50]. Control [HE51, JP59, McL57, Pat53a, PC58, RB50, Shi58a, WD59]. Controlled [RS59]. Conversion [Pra51, Wil59]. Convict [MN52]. Cooper [Scr57]. Coosa [Sco51]. Cope [Zar52]. Coregonus [EB55, Roe58]. Correlation [HG53]. Costs [Kin51]. Cottus [Zar52]. coulteri [EB55]. Counting [Bla57]. County [Coo52a, Coo53, Hun53, Pat53b, She50]. Cover [Vib54]. Crab [Dar59]. Crappie [Hui54, Rei50, Vio53, Wit57]. Creek [KH55, MN52, She50, Ste52, Sum53, WC55]. Creel [MF57, Mur55, TH51]. Cricotopus [Sur59]. Cristivomer [She51]. Culture [RS59, Seg57]. Current [Ken57, McL57, Pra52]. Cutthroat [DHB57, Ech55, Fle52, Irv55, Mil52, Mil54]. Cycle [HE51, Hof59]. Cyprinus [McC52].

Dade [Hun53]. Dakota [Shi58a]. Dale [Par57]. Dam [Bri50, Kat51a]. Daphnia [And50]. Data [Car50b, MKB50]. DDT [Ide57]. Decade [FW57]. December [Ano57j]. Deep [Mof57]. Deep-water [Mof57]. Densities [Bec50]. Deplorable [Gre54]. Depth [Byr52]. Description [GSF53]. Desoto [LP55]. Destruction [LCH59]. Determinant [LTP57]. Determination [AS51, BL50, Ken50]. Determinations [Alv54]. Determining [Ket50, SF52, Syk58]. Development [DO57, Iye50, Lay50, PC58, Vib54, Wol54]. Developments [LP57]. Device [Kin57]. Devices [Moy59]. Diego [Bal58]. Dieldrin [TH57]. Diet [Bon50, ST57]. Different [Bal57]. Differential [Man51]. Direct [McL57, Pra52]. Disc [Bee58]. Disease [SF50, Wol54, Wol57, WD59]. Disease-Resistant [Wol54]. Diseases [All54, Gor54, Hal54b, McL52, REO54, Sni54, Wat54]. Dispersal [Mil57]. Distribution [Byr52]. Distributions [Ket50]. Disturbance [Car58a]. Ditches [HC51]. Dolly [Roo59]. dolomieu [HF57]. Domestic [KH55]. Dorosoma [Mil57]. Downstream [AB52, Wol51]. Drawdowns [Shi58a]. Dried [ST57]. Drought [LCH59]. Dry [Gra57]. duorarum [IM59]. during [Byr52, MJ57]. Dynamics [Tho59].

Early [PS55, Smi57]. East [RM53]. Eastern [Coo52a, Lof58]. Ecological [Ech55, MN52]. Ecology [Hig50, Irv55, JP59, Raw57]. Economic [Sch59]. Economics [Ano55i]. Effect [Gri58, Ide57, Jen57, Roy50, Ruh57, She51, SW55, VAK50, Vib54]. Effecting [CL59]. Effectiveness [CM53, Whi59]. Effects [Bri50, CO57, Eis58, Glu55, KG53, LS54, MN52, MC59, RS59, Saw59, Sey59, Wat57]. Efficiency [Pra51, Pra52, WC55]. Eggs [All58, Eis58, OF57]. Eighth [Ano59g, Ano50v, Ano59c]. Eightieth [Ano51b]. Eighty [Ano52a, Ano57-27, Ano59g, Ano53d, Ano54d, Ano55c, Ano57f, Ano58d, Ano59c]. Eighty-Eighth [Ano59g, Ano59c]. Eighty-Fifth [Ano57-27]. Eighty-First [Ano52a]. Eighty-Fourth [Ano55c]. Eighty-Second [Ano53d]. Eighty-Seventh [Ano58d]. Eighty-Sixth [Ano57f]. Eighty-Third [Ano54d]. Elasticity [Gri59]. Election [Ano51c, Ano54e]. Electric [LP57]. Electrical [Pra55]. Electro [SFK59, WC59]. Electro-Fishing [SFK59, WC59]. Electroplating [Sur59]. Elevation [Pur51]. Elkville [LP55]. Enfield [DB57]. England [Glu55, WJ59]. Equipment [WC59]. Erie [And50, Cle51a, Cle51b, Kin57]. Errata [Ano59d, Ano59e]. Erythrocytes [Kat51b]. Escanaba [Pat53b]. Esox [HC51]. Especially [MJ57]. establishment [LCH59]. Estimate [MT57b]. Estimation [Coo52b]. Etiology [SF50]. European [Hue59]. Evaluation [Sai54]. Expansion [Pow58]. Experiment [HBT53]. Experimental [Ben51, Bon50, Bro51, Mot57, Shi58a, Wol57]. Experiments [Jac50, Swi51]. Exploitation [Coo52a]. Exploration [Pow58, Whi57]. Extinction [Esc57].

f [She50]. Factor [All51]. Factors [Moo57, SF52]. Fall [Ell58, WC55]. Fall-Migrating [Ell58]. False [Spr54b]. Fario [Sig52]. Farm [Bro51, Byr52, Moo57, Zel53]. Farming [TS52]. Fauna [Bri50, CO57, Rug59, Wal51, Woo53]. Federal [Car58b]. Feeding [Ben54, Roo59]. Fence [Bla57]. Fern [DOD59]. Ferricyanide [BL50]. Ferro [BL50]. Ferro- [BL50]. Fertility [BW58]. Fertilization [Bal50b, Lan50]. Fertilized [Bar53, Zel53]. Fertilizer [Jac50]. Field [Dav59]. Fifth [Ano57-27]. Fin [Ger58, Sey59, She51]. Fin-Clipped [Ger58]. Fingerling [Bon50, She50, She51]. Fingerlings [She52]. First [Ano52a, IS52, MJ57]. Fish [All54, BV57, BL50, Car50b, Coo52b, Far54, Ger53, Gor54, GO50, Gri58, Gri54, Hal54b, HW50, HP58, HPT59, Hil50, HG57, HB58, Hue59, Jen57, Joe53, Kat51a, KG53, Kin51, Kin57, Lam59, LCH59, LE53, Man51, McL57, Mil51, Mof57, Moo57, MKB50, Moy50, MC59, Moy59, Pac58, PC58, Pra52, Pra55, RS59, Roa50, Ros51a, RM51, RM53, Ruh57, Sai54, Sco51, Seg57, Shi58b, SFK59, Sni54, Sni58, Ste52, Swi54, TS52, Tan58, TH57, Vla57, Wal51, Wat54, WSV57, Wol51, Wol52, Wol54]. Fish-Impaling [Moy59]. Fish-Planting [HB58]. Fish-Population [MKB50]. Fished [Mil50]. Fisheries [Ano50g, Ano50h, Ano51e, Ano52a, Ano52d, Ano53c, Ano53f, Ano54c, Ano54g, Ano55e, Ano55i, Ano57-27, Ano57e, Ano57i, Ano58c, Ano58f, Ano59c, But50, Car52, Car58b, Deq50, DHB57, Dym57, Hil54, Iye50, Ken57, Kim58, Man50, McK59, MT57a, Nor59, Pow58, Rig58, Sch59, Ste58, Swi59, Tes53, Whi57]. Fishery [Car59, Far57, Gre54, HLN54, HEL51, Lay50, McH59, Ols58, Par57, Par58, Pat53a, Ros51b, Sca54, SBK52, WL58]. Fishes [Bec50, Fun57, HW59, HC51, Idy58, Kna50, Nig54, Pur58]. Fishing [Bal58, BM57, Pfl50, San57, SFK59, Tes53, WC59]. Flathead [MD59]. flavescens [Car50a, PS55, Tha59]. Flood [Pat53a, PC58]. Flood-Control [Pat53a, PC58]. Florida [COHB57, Deq50, DH50, Hui54, Hun53, IS52, PC59, Rei50]. Floridanus [DH50]. Fluctuations [Ben54, SK54]. following [Bec50, RM53]. fontinalis [HE51, She50, SGF50]. Food [Bal57, Bon53, Cle51a, CL59, Dob59, HC51, Hun53, Kna50, Laa51, Pra51, Rei50, Tha59, Wal57, Wil59]. Foods [MN52]. Forest [Ide57, Pac58]. Forests [KH52]. Formation [McF59, Sey59]. Fort [Shi58a]. Four [HPT59, Kat51a]. Fourteen [Roa50]. Fourth [Ano55c]. Francis [Pat53a]. Fresh [Deq50, Rig58]. Fresh-Water [Deq50, Rig58]. Freshwater [HW59, Nig54]. fulvescens [PC55]. Further [Kru50, SF52]. Furunculosis [Fla50, SGF50]. Future [Bus59, Hil54, Pow58, RS59, Ste58]. Fyke [Gri58]. gairdneri [LTP57]. gairdnerii [Bon50]. Game [Kna50, RM53, Tan58]. Game-Fish [RM53]. Gar [Hun53]. Gate [WSV57]. Gear [Pow58, San57]. Genetic [RS59]. Genetics [Gor54]. George [Hui54]. Georgian [Bud57]. Gila [Neu57]. Gill [Ken50, Moy50, CLRC53]. Girard [Neu57]. Gizzard [Jen57, Kut58, Mil57]. Glugea [Hal54a]. Gordy [Ger53]. gracilis [Hof59]. Gravel [Vib54]. Great [Dym57, Far57, Hal54a, Ken57, Mil57, Smi57, Wol52]. Green [Mra52, Hac57, Joe57]. Groups [Bon50]. Growth [All59, AS51, Bal57, Bec50, Bon50, Bon53, BS50, Car50a, Car50b, CLRC53, Cle51b, Gri59, Hil50, HB59, Joe57, Kat51a, KH55, LaP58, LTP57, McC52, MT57b, Neu57, Par54, PC59, PC55, Pur51, Pur58, Roe58, RB50, She52, Sig52, Sne51, SW55, Spr54b, Vib54, Vio53, Wal57, Wil59]. Growth-Control [RB50]. Growths [Nig54]. Grunniens [BS50]. Guiding [BV57]. Gulf [Kna50].

Habitat [Pac58]. Habits [Laa51, Par54, Roo59]. Haddock [Cla59]. Halibut [WJ59]. Hampshire [Hal54a]. Harvest [Idy58, MT57b, RM51]. Harvestability [Man51]. Harvesting [Mil51]. Hatched [FW57]. Hatcheries [Bus59, Gra57, Kin51]. Hatchery [All51, Gre52, Hav59, Mil52, Mil54, Ruc58]. Hatchery-Reared [Hav59, Mil52, Mil54]. Hatching [All58]. Headwaters [LE53, Rup55]. Henry [Irv55]. Herring [SSS59]. hertwigi [Hal54a]. Histories [MJ57]. History [Dar59, Fle52, Hui54, ML55, McC59, PS55, Raw57, Rup59, Zar52]. HOC [Ano52i]. Hollow [Par57]. Holter [Kat51a]. Honorary [Ano51c, Ano54e]. Hoover [Gab50]. Horse [All51]. Houghton [HC51]. Hunt [She50]. Huron [McL52, SE53, Bal50a, Bee58, Bud57, Fry53, Tha59]. Hybrid [Kru50]. Hybridization [DHB57]. Hybrids [BW58]. Hydraulic [GO50]. Hydrocarbon [HPT59].

Ictalurus [AS51, Sne51]. Idaho [Irv55, JP59, Mur55]. Illinois [Ben51, Lar55, LE53, LP55]. Impaling [Moy59]. Impounded [Kim58]. Impoundment [Ros51b, Ruh57, Vio53]. Impoundments [Shi58b, Swi54]. Impressions [GO50]. Improve [RE50]. Improvement [Pac58]. Incidence [Wit57]. Increase [RM53]. Increased [Idy58, RM51]. Increasing [Hel59]. India [Iye50]. Indiana [Ger53]. Indicator [McC52]. Induction [Wol57]. Industry [Gre58, TS52]. Infectious [REO54, SWCP59]. Influence [Ste58]. Injected [Ger58]. Inland [Rup59]. Insecticides [HP58, HPT59]. Insects [Ide57]. Institute [Man50]. Instructions [Ano50g, Ano50h, Ano51e, Ano52d, Ano53f, Ano54g, Ano55e, Ano57i, Ano58f]. Intensive [RM53]. Inter [MJ57]. Inter-Relationships [MJ57]. Interesting [Whi57]. International [Ros51a]. Interracial [DHB57]. interrupta [CLRC53]. Introduction [Cro57, LS54, Whi58]. Invasion [Att59]. Invertebrates [LCH59]. Iowa [CLRC53, McC59, Moo57, RM53]. Island [Bal50a, RM51, FW57]. Israel [TS52].

Jordan [Laa51]. June [Ano51f]. Juvenile [Kut58, Wal57].

Kamloops [LS54]. Kansas [MD59]. Kidney [WD59]. kisutch [All58, All59, DA58]. Known [Phi59]. Kraft [VAK50].

L. [HC51, IS52]. Laboratory [Wil59]. Lac [Raw57]. Lacépède [Pra51, Swi51]. lacustris [AS51, Sne51]. Lake [AB52, Bal50a, Bud57, But50, CLRC53, CL59, DB57, DOD59, EDE53, Esc55, Esc57, Fry53, Ger53, Hac57, HEL51, HBT53, HC51, Irv55, Jen57, Kin57, Kut58, LS54, Lof58, Mar52, Mar57, McC59, MC59, Mra52, MT57b, Neu57, Ols58, Pat53b, Pre57, PC55, RE50, Rei50, Roe58, RM53, Roy50, She51, She52, Spr54b, Tha59, And50, Bee58, Bon53, BG54, Car50a, Cle51a, Cle51b, Cop57, EDE53, EB55, Esc57, HEL51, Hui54, Ken50, McC50, McL52, Mof57, Neu57, Pat53a, Pfl50, PC55, Roa50, RM51, Ros51b, TH51, Wal51, Wit57]. Lakes [Bal50b, Bar53, But50, BM57, Coo52b, Ech55, Far57, JP59, Lam59, Lan50, MJ57, Mar52, MKB50, MF57, PS55, SBK52, SK54, SFK59, Wat57, Whi59, Woo53, Dym57, Far57, Ken57, Mil57, SE53, Smi57]. Lamprey [McL52, Roy50]. Lampreys [AB52, SE53]. Large [Par58, Shi58b]. Largemouth [Ben51, Bro51, COHB57, DH50, Gri59, MT57b, Pra51, Sai58, Swi51, Wil59]. Larvae [Eis58]. Latex [Ger58]. Latex-Injected [Ger58]. Laws [Ano58c]. Legal [Coo53, Moy59]. Legal-Sized [Coo53]. Length [KH55]. Lepomis [Byr52, Swi51]. Lernaean [HW59]. LeSueur [Cle51a, Cle51b, DH50, Rei50]. Levels [Joh57, MC59, SF51, SF52]. Liberalized [Pfl50]. Liberations [CL59]. Life [Dar59, Fle52, Hof59, Hui54, ML55, MJ57, McC59, PS55, Raw57, Rup59, Zar52]. Light [Bee58, Eis58, HE51]. Lime [Wat57]. Limits [Sai58]. Limnological [Smi57]. Linnaeus [HG53, McC52, Sig52]. Lipids [Wil55]. List [Ano50i, Ano50j, Ano51f, Ano52e, Ano53g, Ano54h, Ano55f, Ano57j, Ano58g]. Little [Mur55]. Liver [All51]. Log [Dob59]. Logan [Fle52, Sig52, Zar52]. Longnose [BG54, RE50]. Loss [Vib54]. Lost [RM51]. lota [Cle51a, Cle51b, Cle51a, Cle51b]. Louisa [Woo53]. Louisiana [Dar59, Lam59, Vio53]. Louvers [BV57]. Lower [Kim58, Lay50]. lucius [HC51]. Lymphocystis [Wit57].

macrochirus [Byr52, Swi51]. Macroscopic [CO57]. Maculatus [Rei50]. Maculosa [Cle51a, Cle51b]. Magna [And50]. Maine [DB57, Rup55, Rup59]. Mainstream [Mil51]. Making [GO50]. malma [Roo59]. Manage [McH59]. Management [Bal58, Ben51, BM57, Car58a, Deq50, DHB57, Dym57, Hac57, Hue59, KH52, LW55, MC59, Pac58, Par58, Ros51b, Sai58, Sch59, Shi58b, Tho55b, WL58, Wol52]. Manitoba [Att59, But50, Ken50]. Manitoulin [Bal50a, FW57]. Manuscripts [Ano50g, Ano50h, Ano51e, Ano52d, Ano53f, Ano54g, Ano55e, Ano57i, Ano58f]. Marine [Idy58, Ros51a, Sch59]. marinus [AB52, McL52, Saw59]. Marked [All59, Mra52, She50, She52, SBK52]. Massachusetts’ [Gri58]. Massachussetts [Gri59]. Maturity [Cla59]. Mead [Wal51]. Meander [Roa50]. Measure [Pra52]. Measurements [Hil50]. Meeting [Ano50m, Ano50n, Ano51b, Ano51h, Ano52h, Ano53i, Ano54j, Ano55h, Ano57c, Ano57n, Ano57o, Ano58b, Ano58i, Ano59c, Ano50v, Ano50w, Ano52a, Ano53d, Ano54d, Ano55c, Ano57-27, Ano57f, Ano58d, Ano59g]. Member [Ano54e]. Members [Ano50i, Ano50j, Ano51c, Ano51f, Ano52e, Ano53g, Ano53j, Ano54h, Ano54k, Ano55f, Ano55j, Ano57p, Ano57q, Ano58g, Ano58j]. Membership [Ano57j]. Memoriam [Ano50e, Ano50f, Ano51d, Ano53e, Ano54f, Ano55d, Ano57g, Ano57h, Ano58e, Ano59f]. Menhaden [Kna50]. Mesh [Ken50]. Metabolism [DOD59]. Metals [And50]. Method [Hav59, Hel59, Sne51, Syk58]. Methods [Ell58]. Michigan [AB52, Bal50b, Ben54, Coo52b, Coo52a, Coo53, Cro55, Esc57, Gra57, HBT53, HC51, Mof57, Mra52, Roe58, She50, SE53, Wat57]. Micropterus [DH50, HF57, Pra51, Swi51]. Microsporidan [IM59]. Microsporidian [Hal54a]. Mid [HLN54]. Mid-Century [HLN54]. Midgefly [Sur59]. Midwestern [KG53]. Migrating [Ell58]. Migrations [Sum53]. Mill [VAK50]. Mined [LP55]. Mineral [DOD59]. Minerals [Phi59]. Minnesota [Car50a, MJ57, MKB50, Moy50, MF57, Ols58, PS55, SBK52, SK54]. Minnow [Lar55]. Mississippi [BS50, Nor59]. Missouri [Fun57, Kat51a, Pat53a, PC58, Pur58, Shi58b, Wit57, Zel53]. Mitchell [PS55]. Mitchill [Bal50a, Car50a, Roe58, She50, SBK52, Tha59]. Modification [HE51]. Montana [Ech55, Kat51a, Ste52]. Montmorency [She50]. mordax [Bal50a]. Morone [Bon53, CLRC53]. Mortality [BL50, Mil50, Pra55, She52, Sni58, Tes53]. Movement [AB52, EDE53, Fun57, McL57, SE53, Ste52, WJ59]. Movements [Bud57, Mra52, SBK52]. Moving [Wol51]. Mucus [Saw59]. Mucus-Producing [Saw59]. Mugil [IS52]. Mullet [IS52]. Multiple [Bal58, DWB59, Tan58]. Multiple-Use [DWB59]. Muskegon [Cro55]. Muskellunge [Par59]. Mya [Glu55].

namaycush [Mar52, She51]. National [KH52, Wal51]. Natural [Bal50b, Pre57, Sni58, Tes53]. Nature [Gri54, SWCP59]. Navajo [Neu57]. Navigation [BS50]. near [Esc57, HF57]. Necrosis [SWCP59]. Nets [Gri58, Ken50, Moy50, Whi59]. Netting [Bla57]. Nevada [Wal51]. Niangua [Wit57]. nigro [Rei50]. nigro-Maculatus [Rei50]. Ninth [Ano50w]. Nitrogen [Zel53]. Noc [Roe58]. Nomograph [Hil50]. North [Nig54, SSS59]. Northern [Dob59, Joh57, Roe58, Wil55, Bud57, CO57, JP59, Wat57]. Northwestern [Vio53]. Note [Dob59]. Notes [Ano59i, Ano59j]. Number [Ano59e, Kat51b]. Numerical [Cro55]. Nutrition [Hal54b, Phi59].

Observations [BG54, HW59, HB59, Kru50, Mil50, Wal57]. Oceanography [Idy58]. off [WJ59]. Office [Ano58i]. Officers [Ano50k, Ano50l, Ano50o, Ano51g, Ano51i, Ano52f, Ano52j, Ano53h, Ano53l, Ano54i, Ano55g, Ano55k, Ano57k, Ano57l, Ano57u, Ano57r, Ano58h, Ano58k, Ano59h, Hun50, Spr54a]. Ohio [Pfl50]. Oil [CJ55]. Oklahoma [CF57, Jen57, TH51]. Okoboji [RM53]. Olfactory [HW50]. Oncorhynchus [All58, All59, DA58]. Ontario [Lan50, Mar52, Mar57, McC50, Woo53]. Opening [Ano52g, Ano57m, Gor50, Har53, Hil54, Jam51, Mac50, Pau58, Pri57, Tho55a]. Operators [Ruc58]. Opercular [McC52, McK50, Neu57]. Opinion [WL58]. Optimum [Ken50]. Orange [Rei50]. Oregon [HB58, Sum53]. Organic [HP58]. Organisms [Hun53, VAK50, Wol51]. Organization [Man50]. Origin [Mil57]. Osmerus [Bal50a]. Other [Hun53, Nig54, Wol51]. Otsego [Coo52a, Coo53]. Otter [Kin57]. Our [McH59, Rig58, Pow58]. Owned [BM57]. Oxygen [KPW59, MC59]. Ozarks [Wit57].

Pacific [REO54, Tho59]. Paddlefish [HB59]. Pancreatic [SWCP59]. Panfish [Gri58]. Panguitch [Neu57]. Papers [Ano59i, Ano59j]. Parasite [Hal54a, HW59, IM59]. Parasites [McL52]. Parasitic [All54, SE53]. Parasitic-Phase [SE53]. Park [Lan50, Mar52, Mar57, Woo53]. Particular [McC59, Sai54, Spr54b]. Pathogenic [Gri54]. Paul [CL59, LS54]. Pawhuska [TH51]. Pear [Ste52]. Pelleted [Gra57]. pellucidus [Hof59]. Penaeus [HG53, IM59]. Penetration [Bee58]. Pennsylvania [Ano59c]. Perca [Car50a, PS55, Tha59]. Perch [Car50a, Dob59, Ech55, Gri59, Joe57, MJ57, Mra52, PS55, Tha59]. Personnel [Car59]. Petersen [Dav59]. Peterson [HC51]. Petromyzon [AB52, McL52, Saw59]. Phase [SE53]. PhD [Car52]. Phenols [HW50]. Philadelphia [Ano59c]. Philippine [Man50]. Phosphorus [Bar53, HP58, Zel53]. Photographic [Bon50]. Physico [LP55]. Physico-Chemical [LP55]. Physiology [Phi59]. Pickerel [Gri59]. Pigeon [Ben54, Coo52a, Coo53]. Pike [HC51, Joh57, Wil55]. Pink [IM59]. Place [Tan58]. Places [Ano50m, Ano50n, Ano51h, Ano52h, Ano53i, Ano54j, Ano55h, Ano57n, Ano57o, Ano58i]. Plankton [Kut58]. Planting [HB58]. Plantings [Coo53, She50]. Point [DA58]. Poison [HG57]. Policies [Far54]. Pollutants [HW50]. Polluted [CF57]. Pollution [KG53, KH55]. Pomoxis [Rei50]. Pond [CM53, DB57, Gre52]. Ponds [Bro51, Byr52, Gri58, Kru50, LP55, Moo57, Swi51, Zel53]. Pools [BS50]. Population [Bec50, Ger53, Jen57, KG53, LE53, MKB50, MC59, MT57b, PC58, Pat53b, RE50, SK54, Ste52, Tho59]. Populations [COHB57, Coo52b, Coo52a, Gri58, Lam59, Lof58, Mil50, Mof57, Moo57, Moy50, RM53, Rug59, Ruh57, Sco51, Sni58, Swi54]. Possibilities [Par58]. Possible [Phi59]. Potentials [Rig58]. Pothole [MT57a]. Practices [DWB59]. Prairie [Kut58, MC59]. Predation [Vib54]. Predator [Car58a]. Predator-Prey [Car58a]. Predators [Glu55, She52]. Preparing [Ano50g, Ano50h, Ano51e, Ano52d, Ano53f, Ano54g, Ano55e, Ano57i, Ano58f]. Presence [She52]. Present [Iye50, McK59]. Presentation [Cal54, Gre54]. President [Ano52g, Ano57m, Gor50, Har53, Hil54, Jam51, Mac50, Pau58, Pri57, Tho55a]. Presidents’ [Ano50m, Ano50n, Ano51h, Ano52h, Ano53i, Ano54j, Ano55h, Ano57n, Ano57o, Ano58i]. Presses [GO50]. Prevention [All54]. Prey [Car58a]. Prickley [Ste52]. Private [Ruc58]. Problems [Ano55i, Bel58, Far57, Gre58, Kim58, McC59, Par58, Ruc58, Shi58b, Tho55b]. Procedures [Far54, Sai54, WC59]. Producing [Saw59]. Production [Bel58, BM57, Kin51, PC55, Roa50, SK54]. Productivity [Lar55]. Profiles [Hue59]. Program [Bal58, DOD59, HB58, Lay50]. Progress [Man50]. Propagation [Dym57, Gra57]. Proposed [Iye50]. Prosopium [Laa51]. Pseudoharengus [Mil57]. Public [BM57, WL58]. Puget [All59]. Pulp [VAK50]. Pulsated [McL57]. Punctatus [AS51, Sne51]. Pyatt [LP55]. Pygmy [EB55]. Pyramid [RE50].

Quantitative [Bri50, MF57]. Quebec [Seg57, Vla57].

Raceway [Mot57]. Radiation [Vib54]. Radioactive [Hig50]. Rafinesque [BS50, Byr52, Swi51]. Rainbow [Alv54, Bon50, Coo53, CL59, DO57, LTP57, SW55, ST57]. Rainfall [HG53]. Randall [Shi58a]. Rate [AS51, Bon50, Bon53, BS50, Car50a, Coo52a, Kat51a, LTP57, MT57b, Pur51]. Rates [Bec50, Mil50, Swi51, Tes53, Vio53, Wil59]. Rathbun [Dar59]. Rations [Gra57]. Rays [Sey59]. Re [LCH59]. Re-establishment [LCH59]. Reading [Hel59]. Readings [Bee58]. Reared [Hav59, Mil52, Mil54]. Rearing [Ell58]. Recently [AB52]. Reclaimed [Gri59]. Reclamation [CM53]. Record [Roa50]. Recording [Bon50]. Recreational [Wal51]. Red [Tho59, PS55, SBK52, SK54]. Redear [Ger58]. Redeye [Par54, PC59]. Redrock [Woo53]. Redside [LS54]. Reduced [KPW59]. Reduction [Bec50, RE50, Sni58, SV59]. Reference [McC59, Sai54, SK54, Spr54b]. Refinements [WC59]. Regeneration [Bar53]. Region [Hal54a, PC55]. Registered [Ano53j, Ano54k, Ano55j, Ano57p, Ano57q, Ano58j, Ano59g]. Reintroduction [McC50]. Related [Hue59, Moo57]. Relation [KH55, Kin51, Kna50, Pur51]. Relationship [Bee58, McF59, Rup55]. Relationships [Bar53, Ech55, Hun53, MJ57]. Relative [HPT59]. Release [DA58]. Remarks [Ano52g, Ano57m, Gor50, Har53, Hil54, Jam51, Mac50, Pau58, Pri57, Tho55a]. Removal [Gri58, RM53, She51]. Report [Ano53k, Ano54l, Ano57r, Ano57t, Ano57s, Her54, Hun50, Spr54a, Gab50]. Reports [Ano50o, Ano50t, Ano50r, Ano50p, Ano50u, Ano50s, Ano50q, Ano51i, Ano51k, Ano51j, Ano52i, Ano52j, Ano52l, Ano52k, Ano53l, Ano53n, Ano53m, Ano55k, Ano55n, Ano55m, Ano55l, Ano57c, Ano57u, Ano57x, Ano57w, Ano57v, Ano58b, Ano58k, Ano58l, Ano59h, Hun50, Kru54]. Reproduction [Esc55, HB59, Mar57, Scr57, Shi58a]. Research [Gre54, HLN54, Ken57, Pow58, San57]. Reservoir [DWB59, Pat53a, PC58, Scr57, Shi58a, Tho55b, Par57]. Reservoirs [Bal58, CO57, Mil51, Par58]. Resident [Scr57]. Resistant [Sur59, Wol54]. Resources [McH59, Spr57]. Results [Bro51, DH50, Gre54, Gre52, IS52, RM51, She50]. Return [DA58, Ell58]. Returns [Coo53]. Review [TS52]. Reviews [Ano59a, Ano59b]. Revised [Ano50g, Ano50h, Ano51e, Ano54c, Ano54g, Ano55e, Ano57e, Ano57i, Ano58c, Ano58f]. Richardson [Fle52]. Richland [HF57]. River [AB52, Ben54, Coo52a, Coo53, Cro55, Fle52, Lof58, ML55, Mur55, PC59, Pat53a, Pur58, Rug59, Sco51, Sig52, Zar52, BS50, HF57, Kat51a, Kim58, Lay50, Nor59, OF57, Ros51b, Shi58b]. River-Spawning [Lof58]. Rivers [LaP58, Swi54, Wol52]. Roccus [Dav59, Syk58]. Role [Far57, Phi59, Pow58]. Ronge [Raw57]. Rotary [WSV57]. Rotenone [CM53, CO57, Sai54]. Rough [RM53]. Run [Cro55].

Sac [Wol57]. Saccharomyces [ST57]. Saginaw [Tha59]. Salmo [Bon50, DHB57, Fle52, LTP57, Sig52]. salmoides [DH50, Pra51, Swi51]. Salmon [All58, All59, DA58, Eis58, Ell58, Hav59, HM57, Ide57, Kat51b, McC50, Mur55, OF57, RB50, REO54, Rug59, Sey59, Tho59, Vib54, WC55]. salmonicida [GSF53]. Salmonids [KPW59, Sum53]. Salt [Pur58]. Salvelinus [HE51, Mar52, Roo59, She50, SGF50]. Sampling [Kin57, McC59, Moy50, Sco51]. San [Bal58]. Sand [Sum53]. Santee [Scr57]. sapidissima [Mas52, Wal57]. sapidus [Dar59]. Saskatchewan [Att59, Raw57]. Saugers [Car50a]. saxatilis [Dav59, Syk58]. Scale [Car50b, Hav59, Hel59, Hil50]. Scales [Alv54, GO50, Joe57, Neu57]. Scandinavia [Wol52]. Scientific [Seg57]. Screen [WSV57]. Sculpin [Zar52]. Sea [AB52, McL52, SE53]. Seasonal [Ben54, Wit57]. Secchi [Bee58]. Second [Ano53d]. Secretary [Hun50]. Secretary-Treasurer [Hun50]. Section [Kat51a]. Selection [RS59]. Selective [DO57]. Selectivity [Whi59]. Semiscaber [Zar52]. Seneca [Roy50]. September [Ano59c]. Service [Ano50m, Ano50n, Ano51h, Ano52h, Ano53i, Ano54j, Ano55h, Ano57n, Ano57o, Far54]. Sessional [Ano50a, Ano50b, Ano50p, Ano50q, Ano51a, Ano51j, Ano52b, Ano52k, Ano53a, Ano53m, Ano54a, Ano55a, Ano55l, Ano57a, Ano57b, Ano57c, Ano57v, Ano58a, Ano58b, Kru54]. setiferus [HG53]. Seventh [Ano58d]. Seventy [Ano50v, Ano50w]. Seventy-Eighth [Ano50v]. Seventy-Ninth [Ano50w]. Sewage [KG53]. Sex [Syk58]. Sexual [Cla59, HE51]. Shad [Jen57, Kut58, LaP58, Mas52, Mil57, Wal57]. Shallow [Kut58]. Sheepshead [BS50]. Shell [Glu55]. Shiner [LS54, McC59]. Shocker [LP57]. Shockers [Pra52, Pra55]. Shortened [SV59]. Shorter [Ano59i, Ano59j]. Shrimp [HG53, IM59]. Significance [Hig50]. Silver [All58, All59, DA58, Kat51b]. Simple [Hel59]. Simplified [WSV57]. since [Seg57]. Six [Cop57]. Sixth [Ano57f]. Size [Bon50, Ken50, LTP57, McF59, Sai58]. Size-Recording [Bon50]. Sized [Coo53]. Small [Bri50, HBT53, Jen57, Kru50, Lar55, SFK59]. Smallmouth [FW57, HF57, Wil59]. Smelt [Bal50a, Hal54a, Rup59]. Smith [Car50a]. Society [Ano52a, Ano57-27, Ano59c, Ano53c, Ano54c, Ano57e, Ano58c, Hil54, Ano50g, Ano50h, Ano51e, Ano52d, Ano53f, Ano54g, Ano55e, Ano57i, Ano58f]. Soft [Glu55]. Soft-Shell [Glu55]. Solar [Vib54]. Solution [Far57]. Solutions [BL50]. Some [Bon50, Car50b, Ech55, Eis58, Gre58, Gri58, Gri59, KPW59, LS54, McC59, Moo57, Ruc58, SW55, Whi57, DH50, Mas52, VAK50]. Sound [All59]. Sources [Sca54]. South [Scr57, Shi58a, Bal50a, Fry53]. Southeastern [LP57, Par58]. Southern [Esc55, KH52, LE53, LP55, Mra52]. Southwestern [SV59]. Spaghetti [Dav59]. Spawning [Cro55, HM57, Lof58, Mas52, Pre57, RS59, SSS59, WC55]. Special [Ano50r, Ano50s, Ano50x, Ano50y, Ano53k, Ano54l, Ano55m, Ano57w, Ano57s, Cal54, SK54]. Species [HPT59, Kat51a, SK54]. Spiral [Mot57]. Sport [Bal58, Ols58]. Spottail [McC59]. Spotted [COHB57, Hun53, LE53]. Spraying [Ide57]. Spring [COHB57, Joh57, SV59]. Squawfish [JP59]. St. [Pat53a]. Stabilized [Tes53]. Standing [Ano50t, Ano50u, Ano50z, Ano50-27, Ano51k, Ano51l, Ano52l, Ano52m, Ano53n, Ano53o, Ano54m, Ano55n, Ano55o, Ano57x, Ano57t, Ano57y, Ano57z, Ano58l, Ano58m, Her54]. State [BM57, Gre58, LW55]. State-Owned [BM57]. States [Car58b, DWB59, McK59, Wol52]. Statistics [Cop57, Ger53, Kin51, Ols58]. Status [HEL51, Hil54, Iye50, McK59, Moy59, Wal51]. Steelhead [LW55, ML55]. Stillwater [Gre52]. Stizostedion [Car50a, Raw57, SBK52]. Stock [DO57]. Stocking [Bro51, Gre52, Kru50, Swi51]. Strains [Gre52, Wol54]. Stratification [Byr52]. Stratified [Ket50]. Stream [Bri50, CF57, DB57, Fun57, Hig50, KG53, Lar55, LCH59, LE53, Mil52, Mil54, Ruh57, Rup55]. Streamer [Dav59, Joe53]. Streamer-Type [Joe53]. Streams [Hue59, Ide57, Mas52, McC50, Par59, Ruh57, SV59]. Strength [Joh57]. Strigeoidea [Hof59]. Strip [LP55]. Striped [Dav59, Scr57, Syk58]. Structure [Joe57]. Students [Sca54]. Studies [Car50a, Car50b, Dar59, DH50, HF57, Hof59, Lof58, McC59, SF52]. Study [Mar52]. Sturgeon [PC55]. Sub [Spr57]. Sub-Tropical [Spr57]. Subsampling [Ket50]. Success [Moo57]. Sucker [BG54, RE50]. Suggestions [Sni58]. Sulfamerazine [SF52]. Sulfonamide [Fla50, SGF50, WYS55]. Sulfonamides [SF51, SW55]. Summer [Byr52, MJ57]. Sunfish [COHB57, Ger58, Kru50]. Sunkhaze [Rup55]. Superior [EDE53, EB55, Esc55, HEL51, Lof58]. Supplements [ST57]. Supply [Bal58]. Survey [Car59, DWB59, Nor59]. Surveys [Smi57]. Survival [All58, Ger58, Mil52, Mil54]. Swim [KPW59].

T.V.A. [Whi59]. Tag [Joe53, McK50]. Tagged [Bud57, EDE53, WJ59]. Tagging [DH50, IS52, Vla57]. Tags [Dav59, Vla57]. Tailwater [Par57]. Tamiami [Hun53]. Taxonomy [Ros51a]. Technical [Car59]. Technique [Car58a, Joe53]. Technological [Ste58]. Technology [Ano55i, Man50]. Television [San57]. Temperate [Nig54]. Temperature [COHB57, Glu55, OF57, Pur51, Sey59, SV59]. Temperatures [Bal57]. Ten [HPT59, TH51]. Ten-Year [TH51]. Tennessee [Par59, Ruh57]. Terms [Ano50m, Ano50n, Ano51h, Ano52h, Ano53i, Ano54j, Ano55h, Ano57n, Ano57o, Ano58i]. Test [WD59]. Tests [Bon50, Dav59]. Texas [Bro51, HG53, Kna50]. Texoma [Bon53]. their [Hig50, MJ57]. Theoretical [Tes53]. Therapeutic [WD59]. Therapy [Fla50, Sni54]. there [Man51]. Third [Ano54d]. Thompson [Ech55]. Three [LaP58, Pfl50]. Thresholds [And50]. Time [McF59]. Tissue [SF51, SF52]. Tolerance [OF57]. Torula [Jac50]. Toxaphene [CO57, HG57]. Toxicants [Sai54]. Toxicity [And50, BL50, CJ55, HP58, HPT59, TH57, WYS55]. Trace [DOD59]. Transactions [Ano50g, Ano50h, Ano51e, Ano52d, Ano53f, Ano54g, Ano55e, Ano57i, Ano58f]. Transformed [AB52]. Trap [Wol51]. Trashfish [Gri58]. Trawl [Kin57]. Treasurer [Hun50]. Treatment [All54]. Treatments [SGF50]. Trematoda [Hof59]. Tributary [McC50, Ruh57]. Tropical [Spr57]. Trout [All51, Alv54, Bal57, Bar53, Bel58, Ben54, Bon50, BW58, Bus59, Coo52a, Coo53, CL59, DB57, DO57, DHB57, Ech55, EDE53, Esc55, Esc57, Fla50, Fle52, Fry53, Gra57, Gre52, Gre58, Hac57, HE51, HEL51, Irv55, KH52, LS54, LTP57, LW55, LP57, Lof58, MN52, ML55, Mar52, Mar57, McF59, Mil52, Mil54, MT57a, Mot57, Par57, Phi59, Pre57, Pur51, Roy50, Ruc58, Rup55, She50, She51, She52, Sig52, SF50, SGF50, SF51, SF52, SW55, ST57, SV59, Vib54, WL58, WD59, WYS55]. trutta [Sig52]. Tumors [Nig54]. Tuna [Spr57]. TVA [Mil51]. Two [CO57, MJ57, Mil51, Mur55, Mar52]. Two-Year [Mur55]. Type [Joe53]. Types [Whi59].

U.S. [Far54]. Ulcer [SF50]. Underwater [San57]. Unexploited [Rig58]. Unfertilized [Zel53]. Unfished [Mil50]. Uniformity [Bon50]. United [DWB59, McK59, Wol52]. Unmarked [She52]. upon [Bri50, Car50b, CO57, Gri58, Roy50, Sey59]. Upper [BS50, Nor59]. Upstream [McL57]. Use [BV57, Car50b, Cro55, DWB59, DHB57, HW50, HG57, HM57, Pre57, Bal58, Kru50, Tan58]. Utah [Fle52, Neu57, Sig52, Zar52]. utilis [Jac50, ST57]. Utilization [Kna50, Kut58, Ste58].

Validity [Alv54, Hav59]. Value [Jac50]. Varden [Roo59]. Variation [Rup59, Wit57]. Various [And50, SGF50, SF51, Swi51, Tes53]. Vertebrae [AS51, Sey59]. Vilas [Pat53b]. Virginia [Mas52, Ros51b]. Virus [Wat54]. Vital [Ger53]. vitreum [SBK52, SBK52, Raw57]. Volume [Ano59e].

Walbaum [Dav59, DA58, Roo59, Syk58]. Walleye [MJ57, Ols58, Pat53b, Raw57, SK54]. Walleyes [Cro55, SBK52]. Wappapello [Pat53a]. Washington [HF57, LW55, Rug59]. Wastes [Hig50, Sur59, VAK50]. Water [And50, Bal58, CJ55, Deq50, HW50, Joh57, KPW59, Rig58, Shi58a, SV59, WSV57, Mof57]. Water-Supply [Bal58]. Waters [Dar59, FW57, Gri59, Kim58, KH52, LP57, Moy50, Rup59, Vla57]. Watersheds [Pac58, Tan58]. Wells [CJ55]. Wenatchee [Rug59]. West [IS52]. Western [Hue59, Kin57, Spr57]. White [All51, Bon53, HG53, Wit57]. Whitefish [Bud57, EB55, Laa51, Roe58]. Wild [Coo52a, Gre52, Mil54]. Wildlife [Far54]. williamsoni [Laa51]. Wilson [Mas52]. Winnebago [PC55]. Winter [MN52, MC59]. Winterkill [Bec50]. Wisconsin [Hac57, MT57b, Pat53b, PC55, SV59]. Wood [Jac50]. Woods [Car50a].

Yamag [Hof59]. Year [Fry53, FW57, IS52, Joh57, KH55, Mur55, TH51]. Year-Class [Joh57]. Yearling [CL59]. Years [Cop57, Pfl50, Roa50]. Yeast [Jac50]. Yeasts [ST57]. Yellow [Car50a, CLRC53, Ech55, Gri59, Joe57, MJ57, Mra52, PS55, Raw57, Tha59]. Yellowstone [BG54, Cop57, Laa51]. Yield [CL59]. Yields [FW57, Tes53]. York [Roy50]. Young [Bon53, HC51, OF57, Pra51, Sey59].

References

Applegate:1952:DMR

[AB52] Vernon C. Applegate and Clifford L. Brynildson. Downstream movement of recently transformed sea lampreys, Petromyzon marinus, in the Carp Lake River, Michigan. Transactions of the American Fisheries Society, 81(1):275–290, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allison:1951:HLC

[All51] Leonard N. Allison. Horse liver as the causative factor in white blindness of hatchery brook trout. Transactions of the American Fisheries Society, 80(1):140–147, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allison:1954:APT

[All54] Leonard N. Allison. Advancements in prevention and treatment of parasitic diseases of fish. Transactions of the American Fisheries Society, 83(1):221–228, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1958:STH

[All58] George H. Allen. Survival through hatching of eggs from silver salmon (Oncorhynchus kisutch). Transactions of the American Fisheries Society, 87(1):207–219, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1959:GMS

[All59] George H. Allen. Growth of marked silver salmon (Oncorhynchus kisutch) of the 1950 brood in Puget Sound. Transactions of the American Fisheries Society, 88(4):310–318, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alvord:1954:VAD

[Alv54] William Alvord. Validity of age determinations from scales of brown trout, rainbow trout, and brook trout. Transactions of the American Fisheries Society, 83(1):91–103, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:1950:ATT

[And50] Bertil G. Anderson. The apparent thresholds of toxicity to Daphnia magna for chlorides of various metals when added to Lake Erie water. Transactions of the American Fisheries Society, 78(1):96–113, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:ASCa

[Ano50a] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 78(1):209, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:ASCb

[Ano50b] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 79(1):239, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:CBa

[Ano50c] Anonymous. Concluding business. Transactions of the American Fisheries Society, 78(1):254, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:CBb

[Ano50d] Anonymous. Concluding business. Transactions of the American Fisheries Society, 79(1):307–308, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:Ma

[Ano50e] Anonymous. In memoriam. Transactions of the American Fisheries Society, 78(1):8, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:Mb

[Ano50f] Anonymous. In memoriam. Transactions of the American Fisheries Society, 79(1):8, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:IPMa

[Ano50g] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society: (revised 1949). Transactions of the American Fisheries Society, 78(1):255–260, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:IPMb

[Ano50h] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society: (revised 1950). Transactions of the American Fisheries Society, 79(1):309–314, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:LM

[Ano50i] Anonymous. List of members. Transactions of the American Fisheries Society, 78(1):261–281, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:LMC

[Ano50j] Anonymous. List of members: Compiled April 20, 1950. Transactions of the American Fisheries Society, 79(1):315–338, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:OCa

[Ano50k] Anonymous. Officers and committees, 1947–1948. Transactions of the American Fisheries Society, 78(1):3, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:OCb

[Ano50l] Anonymous. Officers and committees, 1947–1948. Transactions of the American Fisheries Society, 79(1):3, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:PTSa

[Ano50m] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 78(1):6–7, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:PTSb

[Ano50n] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 79(1):6–7, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:RO

[Ano50o] Anonymous. Reports of officers. Transactions of the American Fisheries Society, 79(1):240–254, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:RSCc

[Ano50p] Anonymous. Reports of sessional committees. Transactions of the American Fisheries Society, 78(1):245–253, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:RSCf

[Ano50q] Anonymous. Reports of sessional committees. Transactions of the American Fisheries Society, 79(1):298–306, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:RSCb

[Ano50r] Anonymous. Reports of special committees. Transactions of the American Fisheries Society, 78(1):240–244, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:RSCe

[Ano50s] Anonymous. Reports of special committees. Transactions of the American Fisheries Society, 79(1):278–297, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:RSCa

[Ano50t] Anonymous. Reports of standing committees. Transactions of the American Fisheries Society, 78(1):220–239, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:RSCd

[Ano50u] Anonymous. Reports of standing committees. Transactions of the American Fisheries Society, 79(1):255–277, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:SEA

[Ano50v] Anonymous. Seventy-Eighth Annual Meeting. Transactions of the American Fisheries Society, 78(1):205, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:SNA

[Ano50w] Anonymous. Seventy-Ninth Annual Meeting. Transactions of the American Fisheries Society, 79(1):235, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:SCb

[Ano50x] Anonymous. Special committees. Transactions of the American Fisheries Society, 78(1):5, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:SCd

[Ano50y] Anonymous. Special committees. Transactions of the American Fisheries Society, 79(1):5, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:SCa

[Ano50z] Anonymous. Standing committees, 1947–1948. Transactions of the American Fisheries Society, 78(1):4, January 1950. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1950:SCc

[Ano50-27] Anonymous. Standing committees, 1948–1949. Transactions of the American Fisheries Society, 79(1):4, January 1950. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:ASC

[Ano51a] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 80(1):320, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:EAM

[Ano51b] Anonymous. Eightieth annual meeting. Transactions of the American Fisheries Society, 80(1):315, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:EHM

[Ano51c] Anonymous. Election of honorary members. Transactions of the American Fisheries Society, 80(1):319, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:M

[Ano51d] Anonymous. In memoriam. Transactions of the American Fisheries Society, 80(1):viii, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:IPM

[Ano51e] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society: (revised 1951). Transactions of the American Fisheries Society, 80(1):369–374, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:LMC

[Ano51f] Anonymous. List of members: Compiled June 1, 1951. Transactions of the American Fisheries Society, 80(1):375–398, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:OC

[Ano51g] Anonymous. Officers and committees, 1949–1950. Transactions of the American Fisheries Society, 80(1):iii, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:PTS

[Ano51h] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 80(1):vi–vii, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:RO

[Ano51i] Anonymous. Reports of officers. Transactions of the American Fisheries Society, 80(1):321–327, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:RSCb

[Ano51j] Anonymous. Reports of sessional committees. Transactions of the American Fisheries Society, 80(1):361–368, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:RSCa

[Ano51k] Anonymous. Reports of standing committees. Transactions of the American Fisheries Society, 80(1):328–360, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1951:SC

[Ano51l] Anonymous. Standing committees, 1949–1950. Transactions of the American Fisheries Society, 80(1):iv–v, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:AFS

[Ano52a] Anonymous. American Fisheries Society: Eighty-First Annual Meeting. Transactions of the American Fisheries Society, 81(1): 293, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:ASC

[Ano52b] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 81(1):298, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:CB

[Ano52c] Anonymous. Concluding business. Transactions of the American Fisheries Society, 81(1):352, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:IPM

[Ano52d] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 81(1):355–358, January 1952. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:LM

[Ano52e] Anonymous. List of members. Transactions of the American Fisheries Society, 81(1):359–389, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:OC

[Ano52f] Anonymous. Officers and committees, 1950–1951. Transactions of the American Fisheries Society, 81(1):iii, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:ORP

[Ano52g] Anonymous. Opening remarks of the President. Transactions of the American Fisheries Society, 81(1):294–297, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:PTS

[Ano52h] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 81(1):vi–vii, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:RAH

[Ano52i] Anonymous. Reports of AD HOC committees. Transactions of the American Fisheries Society, 81(1):329–346, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:RO

[Ano52j] Anonymous. Reports of officers. Transactions of the American Fisheries Society, 81(1):299–305, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:RSCb

[Ano52k] Anonymous. Reports of sessional committees. Transactions of the American Fisheries Society, 81(1):347–351, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:RSCa

[Ano52l] Anonymous. Reports of standing committees. Transactions of the American Fisheries Society, 81(1):306–328, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1952:SC

[Ano52m] Anonymous. Standing committees, 1950–1951. Transactions of the American Fisheries Society, 81(1):iv–v, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:ASC

[Ano53a] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 82(1):296, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:CB

[Ano53b] Anonymous. Concluding business. Transactions of the American Fisheries Society, 82(1):338–339, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:CAF

[Ano53c] Anonymous. Constitution of the American Fisheries Society. Transactions of the American Fisheries Society, 82(1):343–355, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:ESA

[Ano53d] Anonymous. Eighty-Second Annual Meeting. Transactions of the American Fisheries Society, 82(1):291, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:M

[Ano53e] Anonymous. In memoriam. Transactions of the American Fisheries Society, 82(1):8, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:IPM

[Ano53f] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society. Transactions of the American Fisheries Society, 82(1):357–360, January 1953. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:LM

[Ano53g] Anonymous. List of members. Transactions of the American Fisheries Society, 82(1):361–391, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:OC

[Ano53h] Anonymous. Officers and committees, 1951–1952. Transactions of the American Fisheries Society, 82(1):3, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:PTS

[Ano53i] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 82(1):6–7, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:RMA

[Ano53j] Anonymous. Registered members in attendance. Transactions of the American Fisheries Society, 82(1):292–293, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:RSCb

[Ano53k] Anonymous. Report of special committee. Transactions of the American Fisheries Society, 82(1):329–330, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:RO

[Ano53l] Anonymous. Reports of officers. Transactions of the American Fisheries Society, 82(1):297–302, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:RSCc

[Ano53m] Anonymous. Reports of sessional committees. Transactions of the American Fisheries Society, 82(1):331–337, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:RSCa

[Ano53n] Anonymous. Reports of standing committees. Transactions of the American Fisheries Society, 82(1):303–328, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1953:SC

[Ano53o] Anonymous. Standing committees, 1951–1952. Transactions of the American Fisheries Society, 82(1):4–5, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:ASC

[Ano54a] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 83(1):360, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:CB

[Ano54b] Anonymous. Concluding business. Transactions of the American Fisheries Society, 83(1):413–414, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:CAF

[Ano54c] Anonymous. Constitution of the American Fisheries Society (revised 1953). Transactions of the American Fisheries Society, 83(1): 417–429, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:ETA

[Ano54d] Anonymous. Eighty-Third Annual Meeting. Transactions of the American Fisheries Society, 83(1):353, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:EHM

[Ano54e] Anonymous. Election of honorary member. Transactions of the American Fisheries Society, 83(1):405, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:M

[Ano54f] Anonymous. In memoriam. Transactions of the American Fisheries Society, 83(1):8, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:IPM

[Ano54g] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society (revised 1954). Transactions of the American Fisheries Society, 83(1):431–434, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:LM

[Ano54h] Anonymous. List of members. Transactions of the American Fisheries Society, 83(1):435–466, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:OC

[Ano54i] Anonymous. Officers and committees, 1952–1953. Transactions of the American Fisheries Society, 83(1):3, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:PTS

[Ano54j] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 83(1):6–7, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:RMA

[Ano54k] Anonymous. Registered members in attendance. Transactions of the American Fisheries Society, 83(1):354–356, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:RSC

[Ano54l] Anonymous. Report of special committees. Transactions of the American Fisheries Society, 83(1):399–404, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1954:SC

[Ano54m] Anonymous. Standing committees, 1952–1953. Transactions of the American Fisheries Society, 83(1):4–5, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:ASC

[Ano55a] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 84(1):321, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:CB

[Ano55b] Anonymous. Concluding business. Transactions of the American Fisheries Society, 84(1):386–387, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:EFA

[Ano55c] Anonymous. Eighty-Fourth Annual Meeting. Transactions of the American Fisheries Society, 84(1):315–316, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:M

[Ano55d] Anonymous. In memoriam. Transactions of the American Fisheries Society, 84(1):8, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:IPM

[Ano55e] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society: (revised 1954). Transactions of the American Fisheries Society, 84(1):389–394, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:LMC

[Ano55f] Anonymous. List of members: Compiled, August, 1955. Transactions of the American Fisheries Society, 84(1):395–426, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:OC

[Ano55g] Anonymous. Officers and committees, 1953–1954. Transactions of the American Fisheries Society, 84(1):3, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:PTS

[Ano55h] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 84(1):6–7, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:PCF

[Ano55i] Anonymous. Problems of the commercial fisheries — conservation, technology, economics. Transactions of the American Fisheries Society, 84(1):297–313, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:RMA

[Ano55j] Anonymous. Registered members in attendance. Transactions of the American Fisheries Society, 84(1):317–318, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:RO

[Ano55k] Anonymous. Reports of officers. Transactions of the American Fisheries Society, 84(1):322–329, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:RSCc

[Ano55l] Anonymous. Reports of sessional committees. Transactions of the American Fisheries Society, 84(1):381–385, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:RSCb

[Ano55m] Anonymous. Reports of special committees. Transactions of the American Fisheries Society, 84(1):372–380, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:RSCa

[Ano55n] Anonymous. Reports of standing committees. Transactions of the American Fisheries Society, 84(1):330–371, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1955:SC

[Ano55o] Anonymous. Standing committees, 1953–1954. Transactions of the American Fisheries Society, 84(1):4–5, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:ASCa

[Ano57a] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 85(1):288, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:ASCb

[Ano57b] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 86(1):447, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:BMR

[Ano57c] Anonymous. Business meeting and reports of sessional committees. Transactions of the American Fisheries Society, 86(1):480–490, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:CB

[Ano57d] Anonymous. Concluding business. Transactions of the American Fisheries Society, 85(1):323, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:CAF

[Ano57e] Anonymous. Constitution of the American Fisheries Society: (revised 1955). Transactions of the American Fisheries Society, 85(1): 325–340, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:ESA

[Ano57f] Anonymous. Eighty-Sixth Annual Meeting. Transactions of the American Fisheries Society, 86(1):433, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:Ma

[Ano57g] Anonymous. In memoriam. Transactions of the American Fisheries Society, 85(1):8, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:Mb

[Ano57h] Anonymous. In memoriam. Transactions of the American Fisheries Society, 86(1):8, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:IPM

[Ano57i] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society: (revised 1954). Transactions of the American Fisheries Society, 86(1):491–496, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:MLC

[Ano57j] Anonymous. Membership list: Compiled, December, 1936. Transactions of the American Fisheries Society, 86(1):497–533, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:OCa

[Ano57k] Anonymous. Officers and committees, 1954–1955. Transactions of the American Fisheries Society, 85(1):3, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:OCb

[Ano57l] Anonymous. Officers and committees, 1955–56. Transactions of the American Fisheries Society, 86(1):3, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:ORP

[Ano57m] Anonymous. Opening remarks of the President. Transactions of the American Fisheries Society, 86(1):436–446, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:PTSa

[Ano57n] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 85(1):6–7, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:PTSb

[Ano57o] Anonymous. Presidents’ terms of service and places of meeting. Transactions of the American Fisheries Society, 86(1):6–7, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:RMAa

[Ano57p] Anonymous. Registered members in attendance. Transactions of the American Fisheries Society, 85(1):283–284, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:RMAb

[Ano57q] Anonymous. Registered members in attendance. Transactions of the American Fisheries Society, 86(1):434–435, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:ROb

[Ano57r] Anonymous. Report of officers. Transactions of the American Fisheries Society, 86(1):448–455, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:RSCe

[Ano57s] Anonymous. Report of special committees. Transactions of the American Fisheries Society, 86(1):476–480, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:RSCd

[Ano57t] Anonymous. Report of standing committees. Transactions of the American Fisheries Society, 86(1):456–476, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:ROa

[Ano57u] Anonymous. Reports of officers. Transactions of the American Fisheries Society, 85(1):289–294, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:RSCc

[Ano57v] Anonymous. Reports of sessional committees. Transactions of the American Fisheries Society, 85(1):314–322, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:RSCb

[Ano57w] Anonymous. Reports of special committees. Transactions of the American Fisheries Society, 85(1):305–313, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:RSCa

[Ano57x] Anonymous. Reports of standing committees. Transactions of the American Fisheries Society, 85(1):295–304, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:SCa

[Ano57y] Anonymous. Standing committees, 1954–1955. Transactions of the American Fisheries Society, 85(1):4–5, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:SCb

[Ano57z] Anonymous. Standing committees, 1955–56. Transactions of the American Fisheries Society, 86(1):4–5, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1957:AFS

[Ano57-27] Anonymous. The American Fisheries Society: Eighty-Fifth Annual Meeting. Transactions of the American Fisheries Society, 85(1): 281–282, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:ASC

[Ano58a] Anonymous. Appointment of sessional committees. Transactions of the American Fisheries Society, 87(1):407, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:BMR

[Ano58b] Anonymous. Business meeting and reports of sessional committees. Transactions of the American Fisheries Society, 87(1):454–460, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:CLA

[Ano58c] Anonymous. Constitution and by-laws of the American Fisheries Society: (revised 1957). Transactions of the American Fisheries Society, 87(1):461–476, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:ESA

[Ano58d] Anonymous. Eighty-Seventh Annual Meeting. Transactions of the American Fisheries Society, 87(1):403, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:M

[Ano58e] Anonymous. In memoriam. Transactions of the American Fisheries Society, 87(1):8, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:IPM

[Ano58f] Anonymous. Instructions for preparing manuscripts for the Transactions of the American Fisheries Society: (revised 1957). Transactions of the American Fisheries Society, 87(1):477–480, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:LMC

[Ano58g] Anonymous. List of members: Compiled April 7, 1958. Transactions of the American Fisheries Society, 87(1):481–523, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:OC

[Ano58h] Anonymous. Officers and committees, 1956–1957. Transactions of the American Fisheries Society, 87(1):3, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:PTO

[Ano58i] Anonymous. Presidents’ terms of Office and places of meeting. Transactions of the American Fisheries Society, 87(1):6–7, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:RMA

[Ano58j] Anonymous. Registered members in attendance. Transactions of the American Fisheries Society, 87(1):404–405, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:RO

[Ano58k] Anonymous. Reports of officers. Transactions of the American Fisheries Society, 87(1):408–415, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:RSC

[Ano58l] Anonymous. Reports of standing committees. Transactions of the American Fisheries Society, 87(1):416–453, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1958:SC

[Ano58m] Anonymous. Standing committees, 1956–57. Transactions of the American Fisheries Society, 87(1):4–5, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1959:BRa

[Ano59a] Anonymous. Book reviews. Transactions of the American Fisheries Society, 88(1):70–71, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1959:BRb

[Ano59b] Anonymous. Book reviews. Transactions of the American Fisheries Society, 88(2):147–148, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1959:EEA

[Ano59c] Anonymous. Eighty-Eighth Annual Meeting of the American Fisheries Society: Philadelphia, Pennsylvania September 8–19, 1958. Transactions of the American Fisheries Society, 88(3):214–224, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1959:E

[Ano59d] Anonymous. Errata. Transactions of the American Fisheries Society, 88(4):358, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1959:EVN

[Ano59e] Anonymous. Errata in volume 88, number 1. Transactions of the American Fisheries Society, 88(3):213, July 1959. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1959:M

[Ano59f] Anonymous. In memoriam — 1958. Transactions of the American Fisheries Society, 88(3):240, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1959:RAE

[Ano59g] Anonymous. Registered Attendance Eighty-Eighth Annual Meeting. Transactions of the American Fisheries Society, 88(3):237–240, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1959:ROC

[Ano59h] Anonymous. Reports of officers and committees. Transactions of the American Fisheries Society, 88(3):225–236, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1959:SPNa

[Ano59i] Anonymous. Shorter papers and notes. Transactions of the American Fisheries Society, 88(1):72–73, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1959:SPNb

[Ano59j] Anonymous. Shorter papers and notes. Transactions of the American Fisheries Society, 88(2):149–153, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Appelget:1951:DAR

[AS51] John Appelget and Lloyd L. Smith, Jr. The determination of age and rate of growth from vertebrae of the channel catfish, Ictalurus lacustris Punctatus. Transactions of the American Fisheries Society, 80(1):119–139, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Atton:1959:IMS

[Att59] F. M. Atton. The invasion of Manitoba and Saskatchewan by carp. Transactions of the American Fisheries Society, 88(3):203–205, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldwin:1950:ASO

[Bal50a] N. S. Baldwin. The American smelt, Osmerus mordax (Mitchill), of South Bay, Manitoulin Island, Lake Huron. Transactions of the American Fisheries Society, 78(1):176–180, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ball:1950:FNL

[Bal50b] Robert C. Ball. Fertilization of natural lakes in Michigan. Transactions of the American Fisheries Society, 78(1):145–155, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldwin:1957:FCG

[Bal57] N. S. Baldwin. Food consumption and growth of brook trout at different temperatures. Transactions of the American Fisheries Society, 86(1):323–328, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ball:1958:MSF

[Bal58] Orville P. Ball. Management of sport fishing in San Diego’s program of multiple use of water-supply reservoirs. Transactions of the American Fisheries Society, 87(1):200–206, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barrett:1953:RBA

[Bar53] Paul H. Barrett. Relationships between alkalinity and adsorption and regeneration of added phosphorus in fertilized trout lakes. Transactions of the American Fisheries Society, 82(1):78–90, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beckman:1950:CGR

[Bec50] William C. Beckman. Changes in growth rates of fishes following reduction in population densities by winterkill. Transactions of the American Fisheries Society, 78(1):82–90, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beeton:1958:RBS

[Bee58] Alfred M. Beeton. Relationship between secchi disc readings and light penetration in Lake Huron. Transactions of the American Fisheries Society, 87(1):73–79, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Belcher:1958:PCT

[Bel58] D. D. Belcher. Problems in commercial trout production. Transactions of the American Fisheries Society, 87(1):368–373, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bennett:1951:ELB

[Ben51] George W. Bennett. Experimental largemouth bass management in Illinois. Transactions of the American Fisheries Society, 80(1): 231–239, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benson:1954:SFF

[Ben54] Norman G. Benson. Seasonal fluctuations in the feeding of brook trout in the Pigeon River, Michigan. Transactions of the American Fisheries Society, 83(1):76–83, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1954:OLS

[BG54] C. J. D. Brown and Richard J. Graham. Observations on the longnose sucker in Yellowstone Lake. Transactions of the American Fisheries Society, 83(1):38–46, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burdick:1950:TFF

[BL50] George Edgar Burdick and Morris Lipschuetz. Toxicity of ferroand ferricyanide solutions to fish, and determination of the cause of mortality. Transactions of the American Fisheries Society, 78(1): 192–202, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blair:1957:CFN

[Bla57] A. A. Blair. Counting fence of netting. Transactions of the American Fisheries Society, 86(1):199–207, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Byrd:1957:PMA

[BM57] I. B. Byrd and D. D. Moss. The production and management of Alabama’s state-owned public fishing lakes. Transactions of the American Fisheries Society, 85(1):208–216, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonham:1950:STE

[Bon50] Kelshaw Bonham. Some tests with experimental groups of fingerling rainbow trout, Salmo gairdnerii, on uniformity and rate of growth, diet, and photographic size-recording. Transactions of the American Fisheries Society, 79(1):94–104, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonn:1953:FGR

[Bon53] Edward W. Bonn. The food and growth rate of young white bass (Morone chrysops) in Lake Texoma. Transactions of the American Fisheries Society, 82(1):213–221, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Briggs:1950:QED

[Bri50] John Carmon Briggs. The quantitative effects of a dam upon the bottom fauna of a small California stream. Transactions of the American Fisheries Society, 78(1):70–81, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1951:RSL

[Bro51] William H. Brown. Results of stocking largemouth black bass and channel catfish in experimental Texas farm ponds. Transactions of the American Fisheries Society, 80(1):210–217, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:1950:ARG

[BS50] Robert Lee Butler and Lloyd L. Smith, Jr. The age and rate of growth of the sheepshead, aplodinotus grunniens Rafinesque, in the Upper Mississippi River navigation pools. Transactions of the American Fisheries Society, 79(1):43–54, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Budd:1957:MTW

[Bud57] John Budd. Movements of tagged whitefish in Northern Lake Huron and Georgian Bay. Transactions of the American Fisheries Society, 86(1):128–134, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buss:1959:TTH

[Bus59] Keen Buss. Trout and trout hatcheries of the future. Transactions of the American Fisheries Society, 88(2):75–80, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:1950:LLF

[But50] George E. Butler. The lakes and lake fisheries of Manitoba. Transactions of the American Fisheries Society, 79(1):18–29, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bates:1957:ULG

[BV57] Daniel W. Bates and Russell Vinsonhaler. Use of louvers for guiding fish. Transactions of the American Fisheries Society, 86(1):38–57, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buss:1958:AFT

[BW58] Keen Buss and James E. Wright. Appearance and fertility of trout hybrids. Transactions of the American Fisheries Society, 87(1):172– 181, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Byrd:1952:DDB

[Byr52] I. B. Byrd. Depth distribution of the bluegill, Lepomis macrochirus Rafinesque, in farm ponds during summer stratification. Transactions of the American Fisheries Society, 81(1):162–170, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Calhoun:1954:PSA

[Cal54] W. T. Calhoun. Presentation of special award. Transactions of the American Fisheries Society, 83(1):369, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1950:GRS

[Car50a] Kenneth D. Carlander. Growth rate studies of saugers, Stizostedion canadense Canadense (Smith) and yellow perch, Perca flavescens (Mitchill) from Lake of the Woods, Minnesota. Transactions of the American Fisheries Society, 79(1):30–42, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1950:SCU

[Car50b] Kenneth D. Carlander. Some considerations in the use of fish growth data based upon scale studies. Transactions of the American Fisheries Society, 79(1):187–194, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1952:WBP

[Car52] Kenneth D. Carlander. What becomes of the PhD’s in fisheries? Transactions of the American Fisheries Society, 81(1):1–3, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1958:DPP

[Car58a] Kenneth D. Carlander. Disturbance of the predator-prey balance as a management technique. Transactions of the American Fisheries Society, 87(1):34–38, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carter:1958:SFA

[Car58b] Bernard T. Carter. The states and federal aid in fisheries. Transactions of the American Fisheries Society, 87(1):398–400, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1959:STF

[Car59] Kenneth D. Carlander. A survey of technical fishery personnel. Transactions of the American Fisheries Society, 88(1):18–22, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1957:BCB

[CF57] Howard P. Clemens and Joe C. Finnell. Biological conditions in a brine-polluted stream in Oklahoma. Transactions of the American Fisheries Society, 85(1):18–27, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1955:TBW

[CJ55] Howard P. Clemens and Woodrow H. Jones. Toxicity of brine water from oil wells. Transactions of the American Fisheries Society, 84(1):97–109, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crossman:1959:YLC

[CL59] E. J. Crossman and P. A. Larkin. Yearling liberations and change of food as effecting rainbow trout yield in Paul Lake, British Columbia. Transactions of the American Fisheries Society, 88(1):36–44, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1959:SMH

[Cla59] John R. Clark. Sexual maturity of haddock. Transactions of the American Fisheries Society, 88(3):212–213, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1951:FBL

[Cle51a] Howard P. Clemens. The food of the burbot Lota lota Maculosa (LeSueur) in Lake Erie. Transactions of the American Fisheries Society, 80(1):56–66, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1951:GBL

[Cle51b] Howard Paul Clemens. The growth of the burbot Lota lota Maculosa (LeSueur) in Lake Erie. Transactions of the American Fisheries Society, 80(1):163–173, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1953:AGC

[CLRC53] Kenneth D. Carlander, William M. Lewis, C. E. Ruhr, and Robert E. Cleary. Abundance, growth, and condition of yellow bass, Morone interrupta Gill, in Clear Lake, Iowa, 1941 to 1951.

Transactions of the American Fisheries Society, 82(1):91–103, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1953:ERP

[CM53] Howard P. Clemens and Mayo Martin. Effectiveness of rotenone in pond reclamation. Transactions of the American Fisheries Society, 82(1):166–177, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cushing:1957:ETR

[CO57] Colbert E. Cushing, Jr. and John R. Olive. Effects of toxaphene and rotenone upon the macroscopic bottom fauna of two Northern Colorado reservoirs. Transactions of the American Fisheries Society, 86(1):294–301, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Caldwell:1957:PSS

[COHB57] David K. Caldwell, Howard T. Odum, Thomas R. Hellier, Jr., and Frederick H. Berry. Populations of spotted sunfish and Florida largemouth bass in a constant-temperature spring. Transactions of the American Fisheries Society, 85(1):120–134, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1952:REW

[Coo52a] Edwin L. Cooper. Rate of exploitation of wild Eastern brook trout and brown trout populations in the Pigeon River, Otsego County, Michigan. Transactions of the American Fisheries Society, 81(1): 224–234, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1952:EFP

[Coo52b] Gerald P. Cooper. Estimation of fish populations in Michigan lakes. Transactions of the American Fisheries Society, 81(1):4–16, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1953:RPL

[Coo53] Edwin L. Cooper. Returns from plantings of legal-sized brook, brown and rainbow trout in the Pigeon River, Otsego County, Michigan. Transactions of the American Fisheries Society, 82(1):265–280, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cope:1957:SYC

[Cop57] Oliver B. Cope. Six years of catch statistics on Yellowstone Lake. Transactions of the American Fisheries Society, 85(1):160–179, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crowe:1955:NAU

[Cro55] Walter R. Crowe. Numerical abundance and use of a spawning run of walleyes in the Muskegon River, Michigan. Transactions of the American Fisheries Society, 84(1):125–136, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crocker:1957:I

[Cro57] Richard Crocker. Introduction. Transactions of the American Fisheries Society, 86(1):383, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:1958:RSS

[DA58] Lauren R. Donaldson and George H. Allen. Return of silver salmon, Oncorhynchus kisutch (Walbaum) to point of release. Transactions of the American Fisheries Society, 87(1):13–22, January 1958. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Darnell:1959:SLH

[Dar59] Rezneat M. Darnell. Studies of the life history of the blue crab (Callinectes sapidus Rathbun) in Louisiana waters. Transactions of the American Fisheries Society, 88(4):294–304, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1959:FTP

[Dav59] William S. Davis. Field tests of Petersen, streamer, and spaghetti tags on striped bass, Roccus saxatilis (Walbaum). Transactions of the American Fisheries Society, 88(4):319–329, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeRoche:1957:LTC

[DB57] Stuart E. DeRoche and Lyndon H. Bond. The lake trout of Cold Stream Pond, Enfield, Maine. Transactions of the American Fisheries Society, 85(1):257–270, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dequine:1950:MFF

[Deq50] John F. Dequine. Management of Florida’s fresh-water fisheries. Transactions of the American Fisheries Society, 78(1):38–41, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dequine:1950:RST

[DH50] John F. Dequine and Charles E. Hall, Jr. Results of some tagging studies of the Florida largemouth bass Micropterus salmoides Floridanus (LeSueur). Transactions of the American Fisheries Society, 79(1):155–166, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:1957:IHC

[DHB57] Lauren R. Donaldson, Donald D. Hansler, and Terry N. Buckridge. Interracial hybridization of cutthroat trout, Salmo clarkii, and its use in fisheries management. Transactions of the American Fisheries Society, 86(1):350–360, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:1957:DRT

[DO57] Lauren R. Donaldson and Paul R. Olson. Development of rainbow trout brood stock by selective breeding. Transactions of the American Fisheries Society, 85(1):93–101, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dobie:1959:NFN

[Dob59] John Dobie. Note on food of northern log perch. Transactions of the American Fisheries Society, 88(3):213, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:1959:FLT

[DOD59] Lauren R. Donaldson, Paul R. Olson, and John R. Donaldson. The fern lake trace mineral metabolism program. Transactions of the American Fisheries Society, 88(1):1–6, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dolan:1959:SMU

[DWB59] Thomas Dolan, Charles B. Wurtz, and Colton H. Bridges. Survey of multiple-use reservoir practices in the United States. Transactions of the American Fisheries Society, 88(1):45–49, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dymond:1957:APM

[Dym57] J. R. Dymond. Artificial propagation in the management of Great Lakes fisheries. Transactions of the American Fisheries Society, 86(1):384–392, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschmeyer:1955:PWC

[EB55] Paul H. Eschmeyer and Reeve M. Bailey. The pygmy whitefish, Coregonus coulteri, in Lake Superior. Transactions of the American Fisheries Society, 84(1):161–199, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Echo:1955:SER

[Ech55] John B. Echo. Some ecological relationships between yellow perch and cutthroat trout in Thompson Lakes, Montana. Transactions of the American Fisheries Society, 84(1):239–248, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschmeyer:1953:MTL

[EDE53] Paul H. Eschmeyer, Russell Daly, and Leo F. Erkkila. The movement of tagged lake trout in Lake Superior, 1950–1952. Transactions of the American Fisheries Society, 82(1):68–77, January 1953. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eisler:1958:SEA

[Eis58] Ronald Eisler. Some effects of artificial light on salmon eggs and larvae. Transactions of the American Fisheries Society, 87(1):151– 162, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1958:RMR

[Ell58] C. H. Ellis. Rearing methods and the return of fall-migrating Chinook salmon. Transactions of the American Fisheries Society, 87(1): 132–138, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschmeyer:1955:RLT

[Esc55] Paul H. Eschmeyer. The reproduction of lake trout in Southern Lake Superior. Transactions of the American Fisheries Society, 84 (1):47–74, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschmeyer:1957:NEL

[Esc57] Paul H. Eschmeyer. The near extinction of lake trout in Lake Michigan. Transactions of the American Fisheries Society, 85(1):102–119, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farley:1954:PPU

[Far54] John L. Farley. Policies and procedures of the U.S. Fish and Wildlife Service. Transactions of the American Fisheries Society, 83(1):13– 19, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Farley:1957:RGL

[Far57] John L. Farley. The role of the Great Lakes Fishery Commission in the solution of Great Lakes problems. Transactions of the American Fisheries Society, 86(1):424–429, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flakas:1950:STF

[Fla50] Kenneth G. Flakas. Sulfonamide therapy of furunculosis in brown trout. Transactions of the American Fisheries Society, 78(1):117– 127, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fleener:1952:LHC

[Fle52] George G. Fleener. Life history of the cutthroat trout, Salmo clarki Richardson, in Logan River, Utah. Transactions of the American Fisheries Society, 81(1):235–248, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fry:1953:YCL

[Fry53] F. E. J. Fry. The 1944 year class of lake trout in South Bay, Lake Huron. Transactions of the American Fisheries Society, 82(1):178– 192, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Funk:1957:MSF

[Fun57] John L. Funk. Movement of stream fishes in Missouri. Transactions of the American Fisheries Society, 85(1):39–57, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fry:1957:YYC

[FW57] F. E. J. Fry and K. E. F. Watt. Yields of year classes of the smallmouth bass hatched in the decade of 1940 in manitoulin Island waters. Transactions of the American Fisheries Society, 85(1):135–143, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gabrielson:1950:CHR

[Gab50] Ira N. Gabrielson. A consideration of the Hoover Report. Transactions of the American Fisheries Society, 79(1):13–17, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gerking:1953:VSF

[Ger53] Shelby D. Gerking. Vital statistics of the fish population of Gordy Lake, Indiana. Transactions of the American Fisheries Society, 82 (1):48–67, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gerking:1958:SFC

[Ger58] Shelby D. Gerking. The survival of fin-clipped and latex-injected redear sunfish. Transactions of the American Fisheries Society, 87(1): 220–228, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glude:1955:ETP

[Glu55] John B. Glude. The effects of temperature and predators on the abundance of the soft-shell clam, Mya arenaria, in New England. Transactions of the American Fisheries Society, 84(1):13–26, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greenbank:1950:HPM

[GO50] John Greenbank and D. John O’Donnell. Hydraulic presses for making impressions of fish scales. Transactions of the American Fisheries Society, 78(1):32–37, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gordon:1950:ORP

[Gor50] Seth Gordon. Opening remarks of the President. Transactions of the American Fisheries Society, 79(1):236–238, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gordon:1954:GFD

[Gor54] Myron Gordon. The genetics of fish diseases. Transactions of the American Fisheries Society, 83(1):229–240, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grassl:1957:PDR

[Gra57] E. F. Grassl. Pelleted dry rations for trout propagation in Michigan hatcheries. Transactions of the American Fisheries Society, 86(1): 307–322, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greene:1952:RSB

[Gre52] C. Willard Greene. Results from stocking brook trout of wild and hatchery strains at Stillwater Pond. Transactions of the American Fisheries Society, 81(1):43–52, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greenbank:1954:DPR

[Gre54] John Greenbank. Deplorable presentation of results of fishery research. Transactions of the American Fisheries Society, 83(1):115– 119, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greene:1958:SSP

[Gre58] A. F. C. Greene. Some state problems associated with the commercial trout industry. Transactions of the American Fisheries Society, 87(1):365–367, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffin:1954:NBP

[Gri54] Philip J. Griffin. The nature of bacteria pathogenic to fish. Transactions of the American Fisheries Society, 83(1):241–253, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grice:1958:ERP

[Gri58] Frank Grice. Effect of removal of panfish and trashfish by fyke nets upon fish populations of some Massachusetts’ ponds. Transactions of the American Fisheries Society, 87(1):108–115, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grice:1959:EGY

[Gri59] Frank Grice. Elasticity of growth of yellow perch, chain pickerel, and largemouth bass in some reclaimed massachussetts waters. Transactions of the American Fisheries Society, 88(4):332–335, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffin:1953:MCD

[GSF53] Philip J. Griffin, S. F. Snieszko, and S. B. Friddle. A more comprehensive description of Bacterium salmonicida. Transactions of the American Fisheries Society, 82(1):129–138, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hacker:1957:BML

[Hac57] Vernon A. Hacker. Biology and management of lake trout in Green Lake, Wisconsin. Transactions of the American Fisheries Society, 86 (1):71–83, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haley:1954:MPG

[Hal54a] A. James Haley. Microsporidian parasite, Glugea hertwigi, in American smelt from the Great Bay Region, New Hampshire. Transactions of the American Fisheries Society, 83(1):84–90, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Halver:1954:FDN

[Hal54b] John E. Halver. Fish diseases and nutrition. Transactions of the American Fisheries Society, 83(1):254–261, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harkness:1953:ORP

[Har53] W. J. K. Harkness. Opening remarks of the President. Transactions of the American Fisheries Society, 82(1):294–295, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:1959:VSM

[Hav59] Keith A. Havey. Validity of the scale method for aging hatcheryreared Atlantic salmon. Transactions of the American Fisheries Society, 88(3):193–196, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horton:1958:OAF

[HB58] Howard F. Horton and Robert L. Borovicka. Oregon’s aerial fishplanting program. Transactions of the American Fisheries Society, 87(1):182–189, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Houser:1959:OGR

[HB59] Alfred Houser and Michael G. Bross. Observations on growth and reproduction of the paddlefish. Transactions of the American Fisheries Society, 88(1):50–52, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hooper:1953:EAC

[HBT53] Frank F. Hooper, Robert C. Ball, and Howard A. Tanner. An experiment in the artificial circulation of a small Michigan lake. Transactions of the American Fisheries Society, 82(1):222–241, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunt:1951:FYP

[HC51] Burton P. Hunt and William F. Carbine. Food of young pike, Esox lucius L., and associated fishes in Peterson’s Ditches, Houghton Lake, Michigan. Transactions of the American Fisheries Society, 80 (1):67–83, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hazard:1951:MSC

[HE51] Thomas P. Hazard and Roland E. Eddy. Modification of the sexual cycle in brook trout (Salvelinus fontinalis) by control of light. Transactions of the American Fisheries Society, 80(1):158–162, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1951:SLT

[HEL51] Ralph Hile, Paul H. Eschmeyer, and George F. Lunger. Status of the lake trout fishery in Lake Superior. Transactions of the American Fisheries Society, 80(1):278–312, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Helm:1959:SMI

[Hel59] William T. Helm. A simple method for increasing contrast in scale reading. Transactions of the American Fisheries Society, 88(4):356, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herrington:1954:RSC

[Her54] William C. Herrington. Report of standing committees. Transactions of the American Fisheries Society, 83(1):370–398, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henderson:1957:SSB

[HF57] Croswell Henderson and Richard F. Foster. Studies of smallmouth black bass (Micropterus dolomieu) in the Columbia River near Richland, Washington. Transactions of the American Fisheries Society, 86(1):112–127, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hildebrand:1953:CRT

[HG53] H. H. Hildebrand and G. Gunter. Correlation of rainfall with the Texas catch of white shrimp, Penaeus setiferus (Linnaeus). Transactions of the American Fisheries Society, 82(1):151–155, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hooper:1957:UTF

[HG57] Frank F. Hooper and Alfred R. Grzenda. The use of Toxaphene as a fish poison. Transactions of the American Fisheries Society, 85(1): 180–190, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Higgins:1950:RWT

[Hig50] Elmer Higgins. Radioactive wastes and their significance in stream ecology. Transactions of the American Fisheries Society, 79(1):217– 232, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1950:NCG

[Hil50] Ralph Hile. A nomograph for the computation of the growth of fish from scale measurements. Transactions of the American Fisheries Society, 78(1):156–162, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hile:1954:ORP

[Hil54] Ralph Hile. Opening remarks of the President: Status and future of the American Fisheries Society. Transactions of the American Fisheries Society, 83(1):357–359, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harkness:1954:FRM

[HLN54] W. J. K. Harkness, Justin W. Leonard, and Paul R. Needham. Fishery research at mid-century. Transactions of the American Fisheries Society, 83(1):212–216, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hourston:1957:UAS

[HM57] W. R. Hourston and D. MacKinnon. Use of an artificial spawning channel by salmon. Transactions of the American Fisheries Society, 86(1):220–230, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoffman:1959:SLC

[Hof59] Glenn L. Hoffman. Studies on the life cycle of Apatemon gracilis pellucidus (Yamag) [Trematoda: Strigeoidea]. Transactions of the American Fisheries Society, 88(2):96–99, April 1959. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henderson:1958:TOP

[HP58] Croswell Henderson and Quentin H. Pickering. Toxicity of organic phosphorus insecticides to fish. Transactions of the American Fisheries Society, 87(1):39–51, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henderson:1959:RTT

[HPT59] C. Henderson, Q. H. Pickering, and C. M. Tarzwell. Relative toxicity of ten chlorinated hydrocarbon insecticides to four species of fish. Transactions of the American Fisheries Society, 88(1):23–32, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huet:1959:PBW

[Hue59] Marcel Huet. Profiles and biology of Western European streams as related to fish management. Transactions of the American Fisheries Society, 88(3):155–163, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huish:1954:LHB

[Hui54] Melvin T. Huish. Life history of the black crappie of Lake George, Florida. Transactions of the American Fisheries Society, 83(1):176– 193, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunter:1950:ROR

[Hun50] Russell P. Hunter. Reports of officers: Report of the Secretary-Treasurer. Transactions of the American Fisheries Society, 78(1): 210–219, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunt:1953:FRB

[Hun53] Burton P. Hunt. Food relationships between Florida spotted gar and other organisms in the Tamiami Canal, Dade County, Florida. Transactions of the American Fisheries Society, 82(1):13–33, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hasler:1950:UFO

[HW50] Arthur D. Hasler and Warren J. Wisby. Use of fish for the olfactory assay of pollutants (phenols) in water. Transactions of the American Fisheries Society, 79(1):64–70, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haley:1959:OLP

[HW59] A. James Haley and Howard E. Winn. Observations on a lernaean parasite of freshwater fishes. Transactions of the American Fisheries Society, 88(2):128–129, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ide:1957:EFS

[Ide57] F. P. Ide. Effect of forest spraying with DDT on aquatic insects of salmon streams. Transactions of the American Fisheries Society, 86(1):208–219, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Idyll:1958:CBO

[Idy58] C. P. Idyll. Contribution of biology and oceanography to increased harvest of marine fishes. Transactions of the American Fisheries Society, 87(1):282–292, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Iversen:1959:NMP

[IM59] Edwin S. Iversen and Raymond B. Manning. A new microsporidan parasite from the pink shrimp (Penaeus duorarum). Transactions of the American Fisheries Society, 88(2):130–132, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Irving:1955:ECT

[Irv55] Robert B. Irving. Ecology of the cutthroat trout in Henry’s Lake, Idaho. Transactions of the American Fisheries Society, 84(1):275– 296, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Idyll:1952:RFY

[IS52] Clarence P. Idyll and John W. Sutton. Results of the first year’s tagging of mullet, Mugil cephalus L., on the West Coast of Florida. Transactions of the American Fisheries Society, 81(1):69–77, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Iyengar:1950:FIP

[Iye50] H. D. R. Iyengar. The fisheries of India — present status and proposed development. Transactions of the American Fisheries Society, 78(1):91–95, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jackson:1950:EVW

[Jac50] C. F. Jackson. Experiments on the value of wood yeast Torula utilis as an aquatic fertilizer. Transactions of the American Fisheries Society, 79(1):105–111, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

James:1951:ORP

[Jam51] Milton C. James. Opening remarks of the President. Transactions of the American Fisheries Society, 80(1):316–318, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jenkins:1957:EGS

[Jen57] Robert M. Jenkins. The effect of gizzard shad on the fish population of a small Oklahoma lake. Transactions of the American Fisheries Society, 85(1):58–74, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Joeris:1953:TAS

[Joe53] Leonard S. Joeris. Technique for the application of a streamer-type fish tag. Transactions of the American Fisheries Society, 82(1): 42–47, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Joeris:1957:SGS

[Joe57] Leonard S. Joeris. Structure and growth of scales of yellow perch of Green Bay. Transactions of the American Fisheries Society, 86(1): 169–194, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1957:NPY

[Joh57] Fritz H. Johnson. Northern pike year-class strength and spring water levels. Transactions of the American Fisheries Society, 86(1):285– 293, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jeppson:1959:ECC

[JP59] Paul W. Jeppson and William S. Platts. Ecology and control of the Columbia squawfish in Northern Idaho lakes. Transactions of the American Fisheries Society, 88(3):197–202, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kathrein:1951:GRF

[Kat51a] Joseph W. Kathrein. Growth rate of four species of fish in a section of the Missouri River between Holter Dam and Cascade, Montana. Transactions of the American Fisheries Society, 80(1):93–98, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Katz:1951:NEB

[Kat51b] Max Katz. The number of erythrocytes in the blood of the silver salmon. Transactions of the American Fisheries Society, 80(1):184– 193, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kennedy:1950:DOS

[Ken50] William A. Kennedy. The determination of optimum size of mesh for gill nets in Lake Manitoba. Transactions of the American Fisheries Society, 79(1):167–179, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kennedy:1957:CFR

[Ken57] W. A. Kennedy. Current fisheries research by Canadians on the Great Lakes. Transactions of the American Fisheries Society, 86(1): 419–423, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ketchen:1950:SSD

[Ket50] K. S. Ketchen. Stratified subsampling for determining age distributions. Transactions of the American Fisheries Society, 79(1):205– 212, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Katz:1953:ESP

[KG53] Max Katz and Arden R. Gaufin. The effects of sewage pollution on the fish population of a Midwestern stream. Transactions of the American Fisheries Society, 82(1):156–165, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

King:1952:MTW

[KH52] Willis King and Ancil Holloway. Management of trout waters in the national forests of the Southern Appalachians. Transactions of the American Fisheries Society, 81(1):53–62, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Katz:1955:LGY

[KH55] Max Katz and W. Charles Howard. The length and growth of 0-year class creek chubs in relation to domestic pollution. Transactions of the American Fisheries Society, 84(1):228–238, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kimsey:1958:FPI

[Kim58] J. B. Kimsey. Fisheries problems in impounded waters of California and the Lower Colorado River. Transactions of the American Fisheries Society, 87(1):319–332, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kingsbury:1951:PSC

[Kin51] Oliver R. Kingsbury. Production statistics and costs in relation to fish hatcheries. Transactions of the American Fisheries Society, 80(1):148–153, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kinney:1957:OTF

[Kin57] Edward C. Kinney. The otter trawl as a fish sampling device in Western Lake Erie. Transactions of the American Fisheries Society, 86(1):58–60, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knapp:1950:MUR

[Kna50] Frank T. Knapp. Menhaden utilization in relation to the conservation of food and game fishes of the Texas Gulf Coast. Transactions of the American Fisheries Society, 79(1):137–144, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Katz:1959:ASS

[KPW59] Max Katz, Austin Pritchard, and Charles E. Warren. Ability of some salmonids and a centrarchid to swim in water of reduced oxygen content. Transactions of the American Fisheries Society, 88 (2):88–95, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krumholz:1950:FOU

[Kru50] Louis A. Krumholz. Further observations on the use of hybrid sunfish in stocking small ponds. Transactions of the American Fisheries Society, 79(1):112–124, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krumholz:1954:RSC

[Kru54] Louis A. Krumholz. Reports of sessional committees. Transactions of the American Fisheries Society, 83(1):406–412, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kutkuhn:1958:UPJ

[Kut58] Joseph H. Kutkuhn. Utilization of plankton by juvenile gizzard shad in a shallow prairie lake. Transactions of the American Fisheries Society, 87(1):80–103, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laakso:1951:FHY

[Laa51] Martin Laakso. Food habits of the Yellowstone whitefish Prosopium williamsoni Cismontanus (Jordan). Transactions of the American Fisheries Society, 80(1):99–109, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lambou:1959:FPB

[Lam59] Victor W. Lambou. Fish populations of backwater lakes in Louisiana. Transactions of the American Fisheries Society, 88(1): 7–15, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Langford:1950:FLA

[Lan50] R. R. Langford. Fertilization of lakes in Algonquin Park, Ontario. Transactions of the American Fisheries Society, 78(1):133–144, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LaPointe:1958:AGA

[LaP58] Donald F. LaPointe. Age and growth of the American shad, from three Atlantic Coast rivers. Transactions of the American Fisheries Society, 87(1):139–150, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larimore:1955:MPS

[Lar55] R. Weldon Larimore. Minnow productivity in a small Illinois stream. Transactions of the American Fisheries Society, 84(1):110–116, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laythe:1950:FDP

[Lay50] Leo L. Laythe. The fishery development program in the Lower Columbia River. Transactions of the American Fisheries Society, 78 (1):42–55, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larimore:1959:DRE

[LCH59] R. Weldon Larimore, William F. Childers, and Carlton Heckrotte. Destruction and re-establishment of stream fish and invertebrates affected by drought. Transactions of the American Fisheries Society, 88(4):261–285, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1953:FPH

[LE53] William M. Lewis and David Elder. The fish population of the headwaters of a spotted bass stream in Southern Illinois. Transactions of the American Fisheries Society, 82(1):193–202, January 1953. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Loftus:1958:SRS

[Lof58] K. H. Loftus. Studies on river-spawning populations of lake trout in Eastern Lake Superior. Transactions of the American Fisheries Society, 87(1):259–277, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1955:PCC

[LP55] William M. Lewis and Charles Peters. Physico-chemical characteristics of ponds in the Pyatt, Desoto, and Elkville strip mined areas of Southern Illinois. Transactions of the American Fisheries Society, 84(1):117–124, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lennon:1957:ESD

[LP57] Robert E. Lennon and Phillip S. Parker. Electric shocker developments on southeastern trout waters. Transactions of the American Fisheries Society, 85(1):234–240, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larkin:1954:SEI

[LS54] P. A. Larkin and S. B. Smith. Some effects of introduction of the redside shiner on the Kamloops trout in Paul Lake, British Columbia. Transactions of the American Fisheries Society, 83(1):161–175, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larkin:1957:SDG

[LTP57] P. A. Larkin, J. G. Terpenning, and R. R. Parker. Size as a determinant of growth rate in rainbow trout Salmo gairdneri. Transactions of the American Fisheries Society, 86(1):84–96, January 1957. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larson:1955:MST

[LW55] Ralph W. Larson and John M. Ward. Management of steelhead trout in the State of Washington. Transactions of the American Fisheries Society, 84(1):261–274, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mackay:1950:ORP

[Mac50] H. H. Mackay. Opening remarks of the President. Transactions of the American Fisheries Society, 78(1):206–208, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mane:1950:PIF

[Man50] Andres M. Mane. The Philippine Institute of Fisheries Technology: Its organization and progress. Transactions of the American Fisheries Society, 79(1):213–216, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manges:1951:THD

[Man51] Daniel E. Manges. Is there a harvestability differential in fish? Transactions of the American Fisheries Society, 80(1):46–49, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:1952:SLT

[Mar52] Nigel V. Martin. A study of the lake trout, Salvelinus namaycush, in two Algonquin Park, Ontario, lakes. Transactions of the American Fisheries Society, 81(1):111–137, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:1957:RLT

[Mar57] Nigel V. Martin. Reproduction of lake trout in Algonquin Park, Ontario. Transactions of the American Fisheries Society, 86(1): 231–244, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Massmann:1952:CSA

[Mas52] William H. Massmann. Characteristics of spawning areas of shad, Alosa sapidissima (Wilson) in some Virginia streams. Transactions of the American Fisheries Society, 81(1):78–93, January 1952. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1959:EMW

[MC59] John B. Moyle and William D. Clothier. Effects of management and winter oxygen levels on the fish population of a prairie lake. Transactions of the American Fisheries Society, 88(3):178–185, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCrimmon:1950:RAS

[McC50] Hugh R. McCrimmon. The reintroduction of Atlantic salmon into tributary streams of Lake Ontario. Transactions of the American Fisheries Society, 78(1):128–132, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McConnell:1952:OBI

[McC52] William J. McConnell. The opercular bone as an indicator of age and growth of the carp, Cyprinus carpio Linnaeus. Transactions of the American Fisheries Society, 81(1):138–149, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCann:1959:LHS

[McC59] James A. McCann. Life history studies of the spottail shiner of Clear Lake, Iowa, with particular reference to some sampling problems. Transactions of the American Fisheries Society, 88(4):336–343, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McFadden:1959:RSA

[McF59] James T. McFadden. Relationship of size and age to time of annulus formation in brook trout. Transactions of the American Fisheries Society, 88(3):176–177, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McHugh:1959:CWM

[McH59] J. L. McHugh. Can we manage our Atlantic coastal fishery resources? Transactions of the American Fisheries Society, 88(2): 105–110, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McKenzie:1950:NCO

[McK50] R. A. McKenzie. A new celluloid opercular tag. Transactions of the American Fisheries Society, 78(1):114–116, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McKernan:1959:PSC

[McK59] Donald L. McKernan. Present status of commercial fisheries in the United States. Transactions of the American Fisheries Society, 88 (3):169–175, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McLain:1952:DPS

[McL52] Alberton L. McLain. Diseases and parasites of the sea lamprey, Petromyzon marinus, in the Lake Huron Basin. Transactions of the American Fisheries Society, 81(1):94–100, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McLain:1957:CUM

[McL57] Alberton L. McLain. The control of the upstream movement of fish with pulsated direct current. Transactions of the American Fisheries Society, 86(1):269–284, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minckley:1959:BFC

[MD59] W. L. Minckley and James E. Deacon. Biology of the flathead catfish in Kansas. Transactions of the American Fisheries Society, 88(4): 344–355, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1957:QCC

[MF57] John B. Moyle and Donald R. Franklin. Quantitative creel census on 12 Minnesota lakes. Transactions of the American Fisheries Society, 85(1):28–38, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1950:OMR

[Mil50] Richard B. Miller. Observations on mortality rates in fished and unfished cisco populations. Transactions of the American Fisheries Society, 79(1):180–186, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1951:FHT

[Mil51] Lawrence F. Miller. Fish harvesting on two TVA mainstream reservoirs. Transactions of the American Fisheries Society, 80(1):2–10, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1952:SHR

[Mil52] Richard B. Miller. Survival of hatchery-reared cutthroat trout in an Alberta stream. Transactions of the American Fisheries Society, 81 (1):35–42, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1954:CSW

[Mil54] Richard B. Miller. Comparative survival of wild and hatchery-reared cutthroat trout in a stream. Transactions of the American Fisheries Society, 83(1):120–130, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1957:ODA

[Mil57] Robert Rush Miller. Origin and dispersal of the alewife, Alosa pseudoharengus, and the gizzard shad, dorosoma cepedianum, in the Great Lakes. Transactions of the American Fisheries Society, 86(1): 97–111, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maloney:1957:LHI

[MJ57] J. E. Maloney and F. H. Johnson. Life histories and interrelationships of walleye and yellow perch, especially during their first summer, in two Minnesota lakes. Transactions of the American Fisheries Society, 85(1):191–202, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1950:FPC

[MKB50] John B. Moyle, Jerome H. Kuehn, and Charles R. Burrows. Fishpopulation and catch data from Minnesota lakes. Transactions of the American Fisheries Society, 78(1):163–175, January 1950. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maher:1955:LHS

[ML55] F. P. Maher and P. A. Larkin. Life history of the steelhead trout of the Chilliwack River, British Columbia. Transactions of the American Fisheries Society, 84(1):27–38, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maciolek:1952:EEW

[MN52] John A. Maciolek and P. R. Needham. Ecological effects of winter conditions on trout and trout foods in Convict Creek, California, 1951. Transactions of the American Fisheries Society, 81(1):202– 217, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moffett:1957:RCD

[Mof57] James W. Moffett. Recent changes in the deep-water fish populations of Lake Michigan. Transactions of the American Fisheries Society, 86(1):393–408, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moorman:1957:SFR

[Moo57] Robert B. Moorman. Some factors related to success of fish populations in Iowa farm ponds. Transactions of the American Fisheries Society, 86(1):361–370, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Motl:1957:ESR

[Mot57] L. F. Motl. Experimental spiral raceway for trout. Transactions of the American Fisheries Society, 86(1):135, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1950:GNS

[Moy50] John B. Moyle. Gill nets for sampling fish populations in Minnesota waters. Transactions of the American Fisheries Society, 79(1):195– 204, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1959:LSF

[Moy59] John B. Moyle. Legal status of fish-impaling devices. Transactions of the American Fisheries Society, 88(3):211–212, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mraz:1952:MYP

[Mra52] Donald Mraz. Movements of yellow perch marked in Southern Green Bay, Lake Michigan, in 1950. Transactions of the American Fisheries Society, 81(1):150–161, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1957:APT

[MT57a] Richard B. Miller and Ronald C. Thomas. Alberta’s ‘pothole’ trout fisheries. Transactions of the American Fisheries Society, 86(1): 261–268, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mraz:1957:AHG

[MT57b] Donald Mraz and C. W. Threinen. Angler’s harvest, growth rate and population estimate of the largemouth bass of Browns Lake, Wisconsin. Transactions of the American Fisheries Society, 85(1): 241–256, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murphy:1955:TYC

[Mur55] Leon W. Murphy. Two-year creel census of the Little Salmon River, Idaho. Transactions of the American Fisheries Society, 84(1):200– 206, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neuhold:1957:AGU

[Neu57] John M. Neuhold. Age and growth of the Utah chub, Gila atraria (Girard), in Panguitch Lake and Navajo Lake, Utah, from scales and opercular bones. Transactions of the American Fisheries Society, 85(1):217–233, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nigrelli:1954:TOA

[Nig54] Ross F. Nigrelli. Tumors and other atypical cell growths in temperate freshwater fishes of North America. Transactions of the American Fisheries Society, 83(1):262–296, January 1954. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nord:1959:FSU

[Nor59] Robert C. Nord. Fisheries survey of Upper Mississippi River. Transactions of the American Fisheries Society, 88(3):210–211, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olson:1957:TTE

[OF57] P. A. Olson and R. F. Foster. Temperature tolerance of eggs and young of Columbia River Chinook salmon. Transactions of the American Fisheries Society, 85(1):203–207, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olson:1958:SWS

[Ols58] Donald E. Olson. Statistics of a walleye sport fishery in a Minnesota lake. Transactions of the American Fisheries Society, 87(1):52–72, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Packer:1958:MFW

[Pac58] Paul E. Packer. Management of forest watersheds and improvement of fish habitat. Transactions of the American Fisheries Society, 87(1):392–397, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsons:1954:GHR

[Par54] John W. Parsons. Growth and habits of the redeye bass. Transactions of the American Fisheries Society, 83(1):202–211, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsons:1957:TFT

[Par57] John W. Parsons. The trout fishery of the tailwater below Dale Hollow Reservoir. Transactions of the American Fisheries Society, 85 (1):75–92, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsons:1958:FMP

[Par58] John W. Parsons. Fishery management problems and possibilities on large southeastern reservoirs. Transactions of the American Fisheries Society, 87(1):333–355, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsons:1959:MTS

[Par59] John W. Parsons. Muskellunge in Tennessee streams. Transactions of the American Fisheries Society, 88(2):136–140, April 1959. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patriarche:1953:FLW

[Pat53a] Mercer H. Patriarche. The fishery in Lake Wappapello, a floodcontrol reservoir on the St. Francis River, Missouri. Transactions of the American Fisheries Society, 82(1):242–254, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patterson:1953:WPE

[Pat53b] Donald L. Patterson. The walleye population in Escanaba Lake, Vilas County, Wisconsin. Transactions of the American Fisheries Society, 82(1):34–41, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pautzke:1958:ORP

[Pau58] Clarence F. Pautzke. Opening remarks of the President. Transactions of the American Fisheries Society, 87(1):406, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Probst:1955:AGP

[PC55] Robert T. Probst and Edwin L. Cooper. Age, growth, and production of the lake sturgeon (Acipenser fulvescens) in the Lake Winnebago Region, Wisconsin. Transactions of the American Fisheries Society, 84(1):207–227, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patriarche:1958:DFP

[PC58] Mercer H. Patriarche and Robert S. Campbell. The development of the fish population in a new flood-control reservoir in Missouri, 1948 to 1954. Transactions of the American Fisheries Society, 87(1): 240–258, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parsons:1959:GRB

[PC59] John W. Parsons and Edward Crittenden. Growth of the redeye bass in Chipola River, Florida. Transactions of the American Fisheries Society, 88(3):191–192, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pflton:1950:TYL

[Pfl50] John Z. Pflton. Three years of liberalized fishing at Lake Alma, Ohio. Transactions of the American Fisheries Society, 78(1):64–69, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1959:KPR

[Phi59] Arthur M. Phillips, Jr. The known and possible role of minerals in trout nutrition and physiology. Transactions of the American Fisheries Society, 88(2):133–135, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powell:1958:REG

[Pow58] Donald E. Powell. The role of exploration and gear research in the future expansion of our commercial fisheries. Transactions of the American Fisheries Society, 87(1):309–315, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prather:1951:EFC

[Pra51] E. E. Prather. Efficiency of food conversion by young largemouth bass Micropterus salmoides (Lacépède). Transactions of the American Fisheries Society, 80(1):154–157, January 1951. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pratt:1952:MEA

[Pra52] Virgil S. Pratt. A measure of the efficiency of alternating and direct current fish shockers. Transactions of the American Fisheries Society, 81(1):63–68, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pratt:1955:FMC

[Pra55] Virgil S. Pratt. Fish mortality caused by electrical shockers. Transactions of the American Fisheries Society, 84(1):93–96, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Prevost:1957:UAN

[Pre57] Gustave Prevost. Use of artificial and natural spawning beds by lake trout. Transactions of the American Fisheries Society, 86(1): 258–260, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pritchard:1957:ORP

[Pri57] A. L. Pritchard. Opening remarks of the President. Transactions of the American Fisheries Society, 85(1):285–287, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pycha:1955:ELH

[PS55] Richard L. Pycha and Lloyd L. Smith, Jr. Early life history of the yellow perch, Perca flavescens (Mitchell), in the Red Lakes, Minnesota. Transactions of the American Fisheries Society, 84(1): 249–260, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Purkett:1951:GRT

[Pur51] Charles A. Purkett, Jr. Growth rate of trout in relation to elevation and temperature. Transactions of the American Fisheries Society, 80(1):251–259, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Purkett:1958:GFS

[Pur58] Charles A. Purkett, Jr. Growth of the fishes in the Salt River, Missouri. Transactions of the American Fisheries Society, 87(1): 116–131, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rawson:1957:LHE

[Raw57] D. S. Rawson. The life history and ecology of the yellow walleye, Stizostedion vitreum, in Lac La Ronge, Saskatchewan. Transactions of the American Fisheries Society, 86(1):15–37, January 1957. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rounsefell:1950:GCC

[RB50] George A. Rounsefell and Lyndon H. Bond. Growth-control charts applied to Atlantic salmon. Transactions of the American Fisheries Society, 78(1):189–191, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rawson:1950:RLS

[RE50] D. S. Rawson and C. A. Elsey. Reduction in the longnose sucker population of Pyramid Lake, Alberta, in an attempt to improve angling. Transactions of the American Fisheries Society, 78(1):13– 31, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reid:1950:FBC

[Rei50] George K. Reid, Jr. Food of the black crappie Pomoxis nigro-Maculatus (LeSueur), in Orange Lake, Florida. Transactions of the American Fisheries Society, 79(1):145–154, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rucker:1954:IDP

[REO54] Robert R. Rucker, Brian J. Earp, and Erling J. Ordal. Infectious diseases of Pacific salmon. Transactions of the American Fisheries Society, 83(1):297–312, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Riggs:1958:UPO

[Rig58] Carl D. Riggs. Unexploited potentials of our fresh-water commercial fisheries. Transactions of the American Fisheries Society, 87(1): 299–308, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:1951:RIF

[RM51] Earl T. Rose and Tom Moen. Results of increased fish harvest in Lost Island Lake. Transactions of the American Fisheries Society, 80(1):50–55, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:1953:IGF

[RM53] Earl T. Rose and Tom Moen. The increase in game-fish populations in East Okoboji Lake, Iowa, following intensive removal of rough fish. Transactions of the American Fisheries Society, 82(1):104– 114, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roach:1950:FYR

[Roa50] Lee S. Roach. Fourteen years record of fish production from Meander Lake. Transactions of the American Fisheries Society, 79(1): 125–136, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roelofs:1958:AGW

[Roe58] Eugene W. Roelofs. Age and growth of whitefish, Coregonus clupeaformis (Mitchill), in Big Bay De Noc and Northern Lake Michigan. Transactions of the American Fisheries Society, 87(1):190–199, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roos:1959:FHD

[Roo59] John F. Roos. Feeding habits of the Dolly Varden, Salvelinus malma (Walbaum), at Chignik, Alaska. Transactions of the American Fisheries Society, 88(4):253–260, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rosa:1951:CIA

[Ros51a] Horacio Rosa, Jr. Considerations for an international approach to taxonomy of marine fish. Transactions of the American Fisheries Society, 80(1):110–118, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rosebery:1951:FMC

[Ros51b] Dean A. Rosebery. Fishery management of Claytor Lake, an impoundment on the New River in Virginia. Transactions of the American Fisheries Society, 80(1):194–209, January 1951. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Royce:1950:ELA

[Roy50] William F. Royce. The effect of lamprey attacks upon lake trout in Seneca Lake, New York. Transactions of the American Fisheries Society, 79(1):71–76, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Riggs:1959:ECS

[RS59] Carl D. Riggs and Kermit E. Sneed. The effects of controlled spawning and genetic selection on the fish culture of the future. Transactions of the American Fisheries Society, 88(1):53–57, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rucker:1958:SPP

[Ruc58] Robert R. Rucker. Some problems of private trout hatchery operators. Transactions of the American Fisheries Society, 87(1):374–379, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruggles:1959:SPB

[Rug59] C. P. Ruggles. Salmon populations and bottom fauna in the Wenatchee River, Washington. Transactions of the American Fisheries Society, 88(3):186–190, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ruhr:1957:ESI

[Ruh57] C. E. Ruhr. Effect of stream impoundment in Tennessee on the fish populations of tributary streams. Transactions of the American Fisheries Society, 86(1):144–157, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rupp:1955:BTR

[Rup55] Robert S. Rupp. Beaver-trout relationship in the headwaters of Sunkhaze Stream, Maine. Transactions of the American Fisheries Society, 84(1):75–85, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rupp:1959:VLH

[Rup59] Robert S. Rupp. Variation in the life history of the American smelt in inland waters of Maine. Transactions of the American Fisheries Society, 88(4):241–252, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saila:1954:BAP

[Sai54] Saul B. Saila. Bio-assay procedures for the evaluation of fish toxicants with particular reference to rotenone. Transactions of the American Fisheries Society, 83(1):104–114, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saila:1958:SLL

[Sai58] Saul B. Saila. Size limits in largemouth black bass management. Transactions of the American Fisheries Society, 87(1):229–239, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sand:1957:AUT

[San57] Reidar F. Sand. Application of underwater television to fishing gear research. Transactions of the American Fisheries Society, 86(1): 158–160, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sawyer:1959:ECC

[Saw59] Philip J. Sawyer. Effects of certain chemicals on mucus-producing cells of Petromyzon marinus. Transactions of the American Fisheries Society, 88(4):305–309, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1952:MMW

[SBK52] Lloyd L. Smith, Jr., Robert L. Butler, and Laurits W. Krefting. Movements of marked walleyes, Stizostedion vitreum Vitreum (Mitchill), in the fishery of the Red Lakes, Minnesota. Transactions of the American Fisheries Society, 81(1):179–196, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scattergood:1954:BSF

[Sca54] Leslie W. Scattergood. Bibliographic sources for fishery students and biologists. Transactions of the American Fisheries Society, 83 (1):20–37, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaefer:1959:BEA

[Sch59] Milner B. Schaefer. Biological and economic aspects of the management of commercial marine fisheries. Transactions of the American Fisheries Society, 88(2):100–104, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scott:1951:SFP

[Sco51] Donald C. Scott. Sampling fish populations in the Coosa River, Alabama. Transactions of the American Fisheries Society, 80(1): 28–40, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scruggs:1957:RRS

[Scr57] George D. Scruggs, Jr. Reproduction of resident striped bass in Santee–Cooper Reservoir, South Carolina. Transactions of the American Fisheries Society, 85(1):144–159, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1953:MPP

[SE53] Bernard R. Smith and Oliver R. Elliott. Movement of parasiticphase sea lampreys in Lakes Huron and Michigan. Transactions of the American Fisheries Society, 82(1):123–128, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seguin:1957:SFC

[Seg57] Louis-Roch Seguin. Scientific fish culture in Quebec since 1945. Transactions of the American Fisheries Society, 86(1):136–143, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seymour:1959:ETU

[Sey59] Allyn Seymour. Effects of temperature upon the formation of vertebrae and fin rays in young Chinook salmon. Transactions of the American Fisheries Society, 88(1):58–69, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snieszko:1950:CEU

[SF50] S. F. Snieszko and S. B. Friddle. A contribution to the etiology of ulcer disease of trout. Transactions of the American Fisheries Society, 78(1):56–63, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snieszko:1951:TLV

[SF51] S. F. Snieszko and S. B. Friddle. Tissue levels of various sulfonamides in trout. Transactions of the American Fisheries Society, 80(1):240–250, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snieszko:1952:FSF

[SF52] S. F. Snieszko and S. B. Friddle. Further studies on factors determining tissue levels of sulfamerazine in trout. Transactions of the American Fisheries Society, 81(1):101–110, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1959:EFS

[SFK59] Lloyd L. Smith, Jr., Donald R. Franklin, and Robert H. Kramer. Electro-fishing for small fish in lakes. Transactions of the American Fisheries Society, 88(2):141–146, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snieszko:1950:VST

[SGF50] S. F. Snieszko, J. S. Gutsell, and S. B. Friddle. Various sulfonamide treatments of furunculosis in brook trout, Salvelinus fontinalis. Transactions of the American Fisheries Society, 78(1):181– 188, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1950:RPM

[She50] David S. Shetter. Results from plantings of marked fingerling brook trout (Salvelinus f. fontinalis Mitchill) in Hunt Creek, Montmorency County, Michigan. Transactions of the American Fisheries Society, 79(1):77–93, January 1950. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1951:EFR

[She51] David S. Shetter. The effect of fin removal on fingerling lake trout (Cristivomer namaycush). Transactions of the American Fisheries Society, 80(1):260–277, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1952:MGM

[She52] David S. Shetter. The mortality and growth of marked and unmarked lake trout fingerlings in the presence of predators. Transactions of the American Fisheries Society, 81(1):17–34, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shields:1958:ECC

[Shi58a] James T. Shields. Experimental control of carp reproduction through water drawdowns in Fort Randall Reservoir, South Dakota. Transactions of the American Fisheries Society, 87(1):23–33, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shields:1958:FMP

[Shi58b] James T. Shields. Fish management problems of large impoundments on the Missouri River. Transactions of the American Fisheries Society, 87(1):356–362, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sigler:1952:AGB

[Sig52] William F. Sigler. Age and growth of the brown trout, Salmo trutta Fario Linnaeus, in Logan River, Utah. Transactions of the American Fisheries Society, 81(1):171–178, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1954:FPA

[SK54] Lloyd L. Smith, Jr. and Laurits W. Krefting. Fluctuations in production and abundance of commercial species in the Red Lakes, Minnesota, with special reference to changes in the walleye population. Transactions of the American Fisheries Society, 83(1):131– 160, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1957:LSG

[Smi57] Stanford H. Smith. Limnological surveys of the Great Lakes — early and recent. Transactions of the American Fisheries Society, 86(1): 409–418, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sneed:1951:MCG

[Sne51] Kermit E. Sneed. A method for calculating the growth of channel catfish, Ictalurus lacustris Punctatus. Transactions of the American Fisheries Society, 80(1):174–183, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snieszko:1954:TBF

[Sni54] S. F. Snieszko. Therapy of bacterial fish diseases. Transactions of the American Fisheries Society, 83(1):313–330, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snieszko:1958:SRN

[Sni58] S. F. Snieszko. Suggestions for reduction of natural mortality in fish populations. Transactions of the American Fisheries Society, 87(1): 380–385, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sprecher:1954:RO

[Spr54a] George E. Sprecher. Report of officers. Transactions of the American Fisheries Society, 83(1):361–368, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sprugel:1954:GBN

[Spr54b] George Sprugel, Jr. Growth of bluegills in a new lake, with particular reference to false annuli. Transactions of the American Fisheries Society, 83(1):58–75, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Springer:1957:TRT

[Spr57] Stewart Springer. Tuna resources of the tropical and sub-tropical Western Atlantic. Transactions of the American Fisheries Society, 85(1):13–17, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scattergood:1959:SNA

[SSS59] Leslie W. Scattergood, Carl J. Sindermann, and Bernard E. Skud. Spawning of North American herring. Transactions of the American Fisheries Society, 88(3):164–168, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stevens:1957:CDY

[ST57] Richard I. Stevens and John R. Truog. Comparison of dried yeasts, Candida utilis and Saccharomyces cerevisiae, as supplements in a diet for rainbow trout. Transactions of the American Fisheries Society, 86(1):161–168, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stefanich:1952:PMF

[Ste52] Frank A. Stefanich. The population and movement of fish in Prickley Pear Creek, Montana. Transactions of the American Fisheries Society, 81(1):260–274, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stern:1958:ITA

[Ste58] Joseph A. Stern. Influence of technological advancements on utilization of fisheries in the future. Transactions of the American Fisheries Society, 87(1):293–298, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sumner:1953:MSS

[Sum53] Francis H. Sumner. Migrations of salmonids in Sand Creek, Oregon. Transactions of the American Fisheries Society, 82(1):139–150, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1959:CBM

[Sur59] Eugene W. Surber. Cricotopus bicinctus, a midgefly resistant to electroplating wastes. Transactions of the American Fisheries Society, 88(2):111–116, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stoeckeler:1959:WTR

[SV59] Joseph H. Stoeckeler and Glenn J. Voskuil. Water temperature reduction in shortened spring channels of Southwestern Wisconsin trout streams. Transactions of the American Fisheries Society, 88(4):286–288, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snieszko:1955:ESS

[SW55] S. F. Snieszko and E. M. Wood. The effect of some sulfonamides on the growth of brook trout, brown trout, and rainbow trout. Transactions of the American Fisheries Society, 84(1):86–92, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snieszko:1959:INP

[SWCP59] S. F. Snieszko, Ken Wolf, Jimmy E. Camper, and Lyle L. Pettijohn. Infectious nature of pancreatic necrosis. Transactions of the American Fisheries Society, 88(4):289–293, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1951:EVR

[Swi51] H. S. Swingle. Experiments with various rates of stocking bluegills, Lepomis macrochirus Rafinesque, and largemouth bass, Micropterus salmoides (Lacépède), in ponds. Transactions of the American Fisheries Society, 80(1):218–230, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1954:FPA

[Swi54] H. S. Swingle. Fish populations in Alabama rivers and impoundments. Transactions of the American Fisheries Society, 83(1):47–57, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1959:CF

[Swi59] H. S. Swingle. Careers in fisheries. Transactions of the American Fisheries Society, 88(1):16–17, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sykes:1958:MDS

[Syk58] James E. Sykes. A method of determining the sex of the striped bass, Roccus saxatilis (Walbaum). Transactions of the American Fisheries Society, 87(1):104–107, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tanner:1958:PGF

[Tan58] Howard A. Tanner. Place of game and fish in multiple use of watersheds. Transactions of the American Fisheries Society, 87(1): 386–391, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tester:1953:TYV

[Tes53] Albert L. Tester. Theoretical yields at various rates of natural and fishing mortality in stabilized fisheries. Transactions of the American Fisheries Society, 82(1):115–122, January 1953. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1951:TYC

[TH51] William H. Thompson and Don Hutson. A ten-year creel census on Lake Pawhuska, Oklahoma. Transactions of the American Fisheries Society, 80(1):11–27, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarzwell:1957:TDF

[TH57] Clarence M. Tarzwell and Croswell Henderson. Toxicity of Dieldrin to fish. Transactions of the American Fisheries Society, 86(1):245– 257, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tharratt:1959:FYP

[Tha59] Robert C. Tharratt. Food of yellow perch, Perca flavescens (Mitchill) in Saginaw Bay, Lake Huron. Transactions of the American Fisheries Society, 88(4):330–331, October 1959. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1955:ORP

[Tho55a] Fred A. Thompson. Opening remarks of the President. Transactions of the American Fisheries Society, 84(1):319–320, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1955:PRM

[Tho55b] William H. Thompson. Problems of reservoir management. Transactions of the American Fisheries Society, 84(1):39–46, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1959:APD

[Tho59] William Francis Thompson. An approach to population dynamics of the Pacific red salmon. Transactions of the American Fisheries Society, 88(3):206–209, July 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tal:1952:RFF

[TS52] S. Tal and M. Shelubsky. Review of the fish farming industry in Israel. Transactions of the American Fisheries Society, 81(1):218– 223, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanHorn:1950:EKP

[VAK50] Willis M. Van Horn, J. B. Anderson, and Max Katz. The effect of Kraft pulp mill wastes on some aquatic organisms. Transactions of the American Fisheries Society, 79(1):55–63, January 1950. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vibert:1954:ESR

[Vib54] Richard Vibert. Effect of solar radiation and of gravel cover on development, growth, and loss by predation in salmon and trout. Transactions of the American Fisheries Society, 83(1):194–201, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Viosca:1953:GRB

[Vio53] Percy Viosca, Jr. Growth rates of black basses and crappie in an impoundment of Northwestern Louisiana. Transactions of the American Fisheries Society, 82(1):255–264, January 1953. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vladykov:1957:FTT

[Vla57] Vadim D. Vladykov. Fish tags and tagging in Quebec waters. Transactions of the American Fisheries Society, 86(1):345–349, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallis:1951:SFF

[Wal51] Orthello L. Wallis. The status of the fish fauna of the Lake Mead National Recreational Area, Arizona–Nevada. Transactions of the American Fisheries Society, 80(1):84–92, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walburg:1957:OFG

[Wal57] Charles H. Walburg. Observations on the food and growth of juvenile American shad, Alosa sapidissima. Transactions of the American Fisheries Society, 86(1):302–306, January 1957. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watson:1954:VDF

[Wat54] Stanely W. Watson. Virus diseases of fish. Transactions of the American Fisheries Society, 83(1):331–341, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:1957:ELA

[Wat57] Thomas F. Waters. The effects of lime application to acid bog lakes in Northern Michigan. Transactions of the American Fisheries Society, 86(1):329–344, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wales:1955:ECS

[WC55] J. H. Wales and Millard Coots. Efficiency of Chinook salmon spawning in Fall Creek, California. Transactions of the American Fisheries Society, 84(1):137–149, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Witt:1959:REP

[WC59] Arthur Witt, Jr. and Robert S. Campbell. Refinements of equipment and procedures in electro-fishing. Transactions of the American Fisheries Society, 88(1):33–35, January 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolf:1959:TTA

[WD59] Ken Wolf and C. E. Dunbar. Test of 34 therapeutic agents for control of kidney disease in trout. Transactions of the American Fisheries Society, 88(2):117–124, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whiteleather:1957:SIA

[Whi57] Richard T. Whiteleather. Some interesting aspects of fisheries exploration — 1956. Transactions of the American Fisheries Society, 86(1):195–198, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whiteleather:1958:I

[Whi58] R. T. Whiteleather. Introduction. Transactions of the American Fisheries Society, 87(1):281, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:1959:SEC

[Whi59] C. E. White, Jr. Selectivity and effectiveness of certain types of commercial nets in the T.V.A. Lakes of Alabama. Transactions of the American Fisheries Society, 88(2):81–87, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilber:1955:LNP

[Wil55] Charles G. Wilber. Lipids in the northern pike. Transactions of the American Fisheries Society, 84(1):150–154, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:1959:FCG

[Wil59] Wells E. Williams. Food conversion and growth rates for largemouth and smallmouth bass in laboratory aquaria. Transactions of the American Fisheries Society, 88(2):125–127, April 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Witt:1957:SVI

[Wit57] Arthur Witt, Jr. Seasonal variation in the incidence of lymphocystis in the white crappie from the Niangua Arm of the Lake of the Ozarks, Missouri. Transactions of the American Fisheries Society, 85(1):271–279, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wise:1959:MTH

[WJ59] John P. Wise and Albert C. Jensen. Movement of tagged halibut off New England. Transactions of the American Fisheries Society, 88(4):357–358, October 1959. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wales:1958:POA

[WL58] J. H. Wales and J. P. Lane. Public opinion analysis in trout fishery management. Transactions of the American Fisheries Society, 87(1):163–171, January 1958. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolf:1951:TCF

[Wol51] Philip Wolf. A trap for the capture of fish and other organisms moving downstream. Transactions of the American Fisheries Society, 80(1):41–45, January 1951. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolf:1952:CAB

[Wol52] Philip Wolf. A comparison of the administrative basis for fish management of rivers in Scandinavia, Great Britain, and the United States. Transactions of the American Fisheries Society, 81(1):197– 201, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolf:1954:DDR

[Wol54] Louis E. Wolf. Development of disease-resistant strains of fish. Transactions of the American Fisheries Society, 83(1):342–349, January 1954. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolf:1957:EIB

[Wol57] Ken Wolf. Experimental induction of blue-sac disease. Transactions of the American Fisheries Society, 86(1):61–70, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wood:1953:BFL

[Woo53] Kenneth G. Wood. The bottom fauna of Louisa and Redrock Lakes, Algonquin Park, Ontario. Transactions of the American Fisheries Society, 82(1):203–212, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whalls:1957:SRF

[WSV57] Marvin J. Whalls, David S. Shetter, and James E. Vondett. A simplified rotary fish screen and an automatic water gate. Transactions of the American Fisheries Society, 86(1):371–380, January 1957. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wood:1955:STB

[WYS55] E. M. Wood, W. T. Yasutake, and S. F. Snieszko. Sulfonamide toxicity in brook trout. Transactions of the American Fisheries Society, 84(1):155–160, January 1955. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zarbock:1952:LHU

[Zar52] William M. Zarbock. Life history of the Utah sculpin, Cottus bairdi Semiscaber (Cope), in Logan River, Utah. Transactions of the American Fisheries Society, 81(1):249–259, January 1952. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zeller:1953:NPC

[Zel53] Howard D. Zeller. Nitrogen and phosphorus concentrations in fertilized and unfertilized farm ponds in Central Missouri. Transactions of the American Fisheries Society, 82(1):281–288, January 1953. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).